### v1

# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

Abstract
Uncrewed Aerial Vehicles (UAVs) are widely deployed across diverse applications due to their mobility and agility. Recent advances in Large Language Models (LLMs) offer a transformative opportunity to enhance UAV intelligence beyond conventional optimization-based and learning-based approaches. By integrating LLMs into UAV systems, advanced environmental understanding, swarm coordination, mobility optimization, and high-level task reasoning can be achieved, thereby allowing more adaptive and context-aware aerial operations. This survey systematically explores the intersection of LLMs and UAV technologies and proposes a unified framework that consolidates existing architectures, methodologies, and applications for UAVs. We first present a structured taxonomy of LLM adaptation techniques for UAVs, including pretraining, fine-tuning, Retrieval-Augmented Generation (RAG), and prompt engineering, along with key reasoning capabilities such as Chain-of-Thought (CoT) and In-Context Learning (ICL). We then examine LLM-assisted UAV communications and operations, covering navigation, mission planning, swarm control, safety, autonomy, and network management. After that, the survey further discusses Multimodal LLMs (MLLMs) for human–swarm interaction, perception-driven navigation, and collaborative control. Finally, we address ethical considerations, including bias, transparency, accountability, and Human-in-the-Loop (HITL) strategies, and outline future research directions. Overall, this work positions LLM-assisted UAVs as a foundation for intelligent and adaptive aerial systems.
Index Terms:
Uncrewed Aerial Vehicles, Large Language Models, 6G Networks, Retrieval-Augmented Generation, Prompt Engineering, In-Context Learning, Chain-of-Thought.
I Introduction
Uncrewed Aerial Vehicles (UAVs) have become a salient enabling technology across a wide range of applications, including environmental monitoring [74], public safety [105], transportation [85], agriculture [92], and infrastructure inspection [4]. For example, UAVs can track environmental changes, support emergency response during disasters, assess crop health, deliver parcels, and assist with traffic management. Due to their flexibility and mobility, UAVs provide efficient and cost-effective solutions in situations where conventional ground-based systems are limited or infeasible [56].
From a communication perspective, UAVs are increasingly important for enhancing wireless networks. They can serve as flying base stations to expand network coverage, especially in areas with sparse or damaged infrastructure [40, 3]. UAVs can also operate as relays [64], [95], collect data from ground devices [63], and provide edge computing capabilities in the air [65], reducing communication latency. On the other hand, UAVs may operate as cellular users, requiring reliable, low-latency connectivity for safe navigation, control, and sensing [6]. When multiple UAVs cooperate, they can form Flying Ad hoc Networks (FANETs) that self-organize and provide either extended range [94] or communication services over large geographical areas or in environments without pre-existing infrastructure. These capabilities make UAVs a critical component of wireless backhauling, smart city applications, and future mobile communication systems [87].
Advancements in cost efficiency, operational flexibility, and dynamic deployment capabilities are driving the rapid expansion of the UAV ecosystem. Modern UAV platforms are increasingly equipped with high-resolution cameras and precision sensors, enabling large-scale and cost-effective data acquisition. As a result, UAVs have become highly valuable for numerous civilian applications across diverse industries [44]. The global UAV market was valued at USD 36.41 billion in 2024 and is projected to reach USD 125.91 billion by 2032, with a Compound Annual Growth Rate (CAGR) of 17.3% [48]. This rapid market growth underscores the growing demand for intelligent UAV-assisted solutions.
In parallel with these developments, Artificial Intelligence (AI) has become a salient enabler for enhancing the autonomy and efficiency of UAV-assisted communication systems. AI techniques are widely used to optimize UAV mobility, resource allocation, and scheduling, allowing UAVs to adapt dynamically to both aerial and ground operations. These approaches support large-scale network management by enabling efficient and automated data collection. The adoption of AI is primarily driven by the inherent complexity and highly dynamic nature of UAV operating environments, where accurate analytical modeling is often infeasible. Additionally, UAV mobility creates high-dimensional state and action spaces, whereas limited sensing capabilities and intermittent connectivity result in incomplete or outdated information about ground devices, including battery levels, energy availability, and channel conditions. These challenges make data-driven and learning-based solutions particularly well-suited to UAV network optimization [48].
Recently, AI has advanced significantly with the emergence of Large Language Models (LLMs), which provide enhanced capabilities in language understanding, generation, and reasoning. Building on decades of progress in natural language processing and Machine Learning (ML), LLMs have become a major driver of recent developments in AI-assisted systems. Their ability to interpret complex instructions and perform multi-step reasoning has generated substantial interest across a wide range of application domains. In UAV systems, LLMs offer new opportunities to automate complex decision-making processes, enhance situational awareness, and support intelligent network operations [53]. However, deploying LLM-assisted solutions in safety- and mission-critical environments introduces significant ethical and operational challenges, including data privacy and security, bias and fairness, and accountability and transparency. Addressing these concerns is essential to ensure the trustworthiness and reliable integration of LLMs.
Moreover, while state-of-the-art LLMs primarily operate on textual data, UAV-assisted networks inherently generate heterogeneous information, including wireless signals, visual sensing data, and contextual metadata. This modality mismatch has motivated the development of Multimodal LLMs (MLLMs) and Vision-Language Models (VLMs) that can jointly process and reason over diverse data sources [43]. This multimodal intelligence is particularly relevant to UAVs, where effective decision-making often depends on integrating communication, sensing, and environmental information.
I-A Why Combining UAVs and LLMs Matters
The integration of UAVs with LLMs introduces a new level of intelligence that conventional control and optimization methods struggle to achieve. Optimization has long been central to UAV communications, such as resource allocation and trajectory design. These problems typically require selecting optimal solutions from discrete decision spaces to maximize objectives, such as throughput, under strict operational constraints. Conventional optimization techniques, such as convex optimization and metaheuristic algorithms, lack sufficient adaptability in highly dynamic UAV environments. Although ML and Deep Reinforcement Learning (DRL) techniques offer greater flexibility, they often exhibit limited generalization and a persistent simulation-to-reality gap, hindering their reliable deployment in real-world UAV scenarios.
Compared with conventional ML-based optimization methods, LLMs introduce a fundamentally different paradigm centered on semantic reasoning rather than parameter tuning alone. While traditional ML and DRL approaches optimize numerical objectives through repeated interaction with simulated environments, LLMs operate at a higher abstraction level, enabling structured reasoning over heterogeneous inputs and mission-level objectives. This capability allows LLMs to function not merely as optimizers, but as cognitive coordinators within UAV systems. Specifically, in UAV communications and networking scenarios, LLMs can interpret network logs, topology information, channel conditions, and mission constraints in natural language form, transforming raw telemetry into structured situational awareness. Instead of solving isolated optimization subproblems (e.g., power allocation or trajectory refinement), LLMs can synthesize cross-layer information to guide adaptive spectrum management, cooperative relay strategies, and swarm-level resource coordination. By reasoning over operational rules, safety constraints, and mission priorities, LLMs can generate context-aware decision strategies that complement lower-level control modules.
Beyond communication optimization, LLMs enhance UAV autonomy by bridging perception, planning, and execution. Through multimodal integration, LLMs can align sensor data, map representations, and textual mission descriptions into a unified reasoning framework. This enables dynamic task decomposition, contingency planning, and collaborative swarm coordination under uncertain or partially observable conditions. Rather than relying solely on fixed reward functions, LLM-driven systems can adjust strategies based on evolving mission semantics and operator intent. In addition, this reduces the dependence on handcrafted rule sets and specialized pipelines and mitigates the need for multiple task-specific models. By unifying diverse data modalities—natural-language commands, telemetry streams, network metrics, and environmental semantics—LLMs provide a coherent operational layer that supports complex, multi-objective UAV missions.
A defining feature of LLMs is In-Context Learning (ICL), which enables real-time adaptation to user requirements through feedback alone, without the computational overhead of retraining. This adaptability, combined with inherent scalability, allows LLMs to handle the combinatorial complexity of multi-UAV optimization by decomposing high-level, multi-objective missions into coherent sub-tasks and executable strategies. Furthermore, the rich semantic knowledge embedded in LLMs supports decision-making that closely aligns with human intuition and established operational doctrines. These reasoning capabilities also improve transparency and trust by enabling UAV systems to generate natural-language explanations of their actions. This explainability enables more intuitive communication with human operators, reduces cognitive and operational burdens, and allows non-expert users to control UAVs effectively. Ongoing advances in communication technologies, such as 6G and satellite internet, provide low-latency, high-bandwidth connectivity required to support real-time LLM-assisted UAV operations, thereby accelerating the transition toward intelligent, autonomous aerial systems.
In the following, we present a unified abstraction at the system-level to explain how LLMs interact holistically with UAV communication, control, and decision-making. This is a macro-level methodological framework for understanding LLM–UAV integration that is organized into three layers:
  • At the highest layer, the cognitive and decision-making layer assigns LLMs the role of meta-controller, responsible for high-level reasoning and mission orchestration. Rather than operating on raw sensor data, the LLM processes symbolic and semantic representations of mission goals, environmental context, and system constraints. This layer translates UAV intent and abstract objectives into structured task descriptions, enabling mission planning, task decomposition, constraint reasoning, and adaptive replanning in dynamic environments.
  • The perception layer mediates between cognition and execution by converting multimodal sensor inputs into semantically meaningful descriptions. Data from vision, LiDAR, radar, and spectrum sensors are fused into higher-level concepts such as objects, activities, behaviors, and inferred intent. LLMs, often combined with MLLMs, provide contextual interpretation, allowing the UAV to understand not only what is observed but also how it relates to mission objectives.
  • The control and communication layer, at the bottom, translate high-level plans and semantic decisions into executable actions, such as control commands, flight trajectories, and communication policies. Execution outcomes and system states are abstracted and relayed to higher layers, forming a closed-loop learning and adaptation process.
This layered abstraction unifies the diverse LLM-assisted methods surveyed in this paper by framing LLMs as cross-layer cognitive integrators rather than isolated components. It clarifies how LLMs complement traditional UAV subsystems, enabling scalable autonomy, semantic reasoning, and adaptive decision-making in communication, control, and mission planning.
I-B Main Contributions
This paper presents a comprehensive investigation of LLM-UAV interactions. It systematically examines how LLM-assisted techniques, such as Retrieval-Augmented Generation (RAG) and prompt engineering, are fundamentally transforming UAV operations by enhancing decision-making, enabling adaptive path planning, and supporting real-time human–UAV interaction. In addition, we explore the paradigm shift from text-only LLMs to MLLMs, analyzing their architectures, training methodologies, and emerging applications in vision–language navigation, swarm coordination, and high-level mission planning. After that, beyond technical advancements, we critically assess the ethical implications of integrating LLMs into UAV operation and communications, including privacy and data security, bias and fairness, transparency and accountability, and environmental sustainability. Finally, we present an LLM-aided UAV case study and identify the future research directions. Overall, this work serves as a comprehensive technical survey and tutorial, as well as an essential ethical guide for the evolving landscape of LLM-assisted autonomous UAV systems.
The main contributions of this survey are the following:
  • Comprehensive synthesis and unified framework that systematically maps the convergence of LLMs and UAV systems across operational, communication, and ethical dimensions. By consolidating existing fragmented research, it provides a unified framework and a clear overview of existing architectures, methodologies, and real-world applications. In addition, we introduce a structured taxonomy of LLM approaches explicitly tailored to UAV challenges, encompassing pretraining, fine-tuning, RAG, and prompt engineering. This taxonomy clarifies the current research landscape and highlights promising directions for future development at the intersection of LLMs and UAVs.
  • In-depth discussion of MLLMs as a transformative enabler for next-generation UAV systems, detailing their core architectures, training paradigms, and emerging capabilities such as Multimodal Chain of Thought (MCoT) reasoning and Multimodal ICL (M-ICL). We further demonstrate how MLLMs enable vision–language navigation and facilitate coordinated swarm control, underscoring their growing importance in complex perception, reasoning, and decision-making tasks for autonomous UAVs.
  • Rigorous analysis of the ethical challenges associated with deploying LLM-assisted UAV systems (LAUS), including risks related to bias and fairness, transparency and accountability in autonomous decision-making, and the environmental impact of LLMs. Beyond identifying these issues, the survey overviews related mitigation strategies, emphasizing LLM observability for systematic monitoring and debugging. Together, these recommendations establish a practical foundation for building trustworthiness, responsibility, and resilience in LLM-assisted UAV communications.
I-C Survey of Surveys
Table I summarizes key surveys on integrating LLMs with UAVs. Early comprehensive research by Javaid et al. [49] examines how LLMs can enhance UAV autonomy, intelligence, and operational efficiency. Their survey reviews a wide range of notable LLM architectures, including Bidirectional Encoder Representations from Transformers (BERT) [21], Generative Pretrained Transformers (GPT), T5, XLNet, and evaluates their applicability to various UAV tasks such as natural language-based control, real-time data analysis, and mission planning. This work lays the foundation for understanding LLM-aided UAV’s technical potential.
Table I: A summary of existing surveys on the integration of LLMs and MLLMs in UAV Operation and Communication.
Surveys Focus & Scope Key Findings / Application Categories
[49] A broad overview of integrating various LLMs (BERT, GPT, T5, etc.) with UAVs to enhance autonomous systems. Applications in natural language control, real-time data analysis, and mission planning for domains like surveillance and disaster response.
[16] Analyzes 17 papers to map the research landscape (2021–2024). Applications: UAV Optimization (35%), Adaptation (40%), Communication (25%). Models: GPT most used (58%). Highlights challenges including hardware limitations, latency, security, and ethical alignment.
[130] Focuses on revolutionizing UAV-assisted wireless communication, information perception, and autonomous control. Enhances aerial networking, edge computing, multimodal data fusion, scene understanding, and path planning.
[54] Comprehensively surveys Transformer-based models (ViTs, STTs, etc.) and LLMs for UAV systems. Categorizes architectures for object detection, tracking, navigation, and mission planning. Highlights LLMs for high-level reasoning and interaction.
[14] Analyzes 74 papers and 56 projects across 9 core tasks in the UAV workflow. Academia: Broad, theoretical focus (e.g., multi-UAV planning). Industry: Practical focus on flight control, single-UAV planning, and human–machine interaction.
[62] A systematic review of LLM and MLLM integration in autonomous driving systems. Explores applications in perception, prediction, planning, decision-making, multitasking processing, and HMI. Discusses multimodal fusion, prompt engineering, and distillation, outlining challenges and future directions.
[18] Systematic investigation of MLLMs for autonomous driving and mapping systems. Reviews MLLM tools, datasets, and benchmarks; summarizes results from the LLVM-AD workshop; identifies major challenges for future adoption in driving systems.
This survey LLMs and MLLMs that enhance UAV autonomy in navigation, communication, and swarm coordination, using techniques such as fine-tuning, RAG, and prompt engineering. The survey addresses ethical issues, including bias and transparency, and recommends HITL oversight to ensure safety. Practical and future research directions are provided. It positions LLM-assisted UAVs as a transformative, intelligent aerial framework. LLMs enhance UAV navigation, path planning, and real-time swarm coordination, surpassing traditional optimization methods. In communications, they optimize spectrum management, resource allocation, and the integration of 5G/6G networks. MLLMs enable multimodal tasks like search and rescue by fusing visual, sensor, and language data. The technology also addresses ethical deployment through safety verification and human oversight frameworks.
Building on this foundation, Cidjeu et al. [16] present a structured mapping of research published between 2021 and 2024, categorizing 17 representative studies into UAV optimization, adaptation, and communication. Their analysis reveals that GPT-based models currently dominate integration efforts, highlighting ongoing challenges related to onboard computational constraints, latency, security, and system robustness. Similarly, Yumeng et al. [130] examine the transformative role of LLMs in UAV-assisted wireless communications, multimodal perception, and autonomous control. Their survey demonstrates how LLMs support edge computing, aerial networking, and trajectory planning through enhanced reasoning and language understanding. They also emphasize the importance of addressing issues such as model adaptation, reliability, and real-time deployment in dynamic environments.
Complementing these LLM-assisted studies, Kheddar et al. [54] broaden the scope by surveying transformer-based models for UAV applications, ranging from hybrid Convolutional Neural Network (CNN)-Transformer architectures to vision and spatio-temporal transformers. Their taxonomy includes contributions in object detection, tracking, and mission planning, and highlights the growing importance of LLMs for high-level reasoning and human–UAV interaction. However, they also identify critical challenges, including data scarcity, multimodal sensor fusion, and the difficulty of meeting real-time operational requirements.
A more extensive empirical study by Chen et al. [14] analyzed 74 academic publications and 56 industrial projects to identify nine core LLM-assisted tasks spanning the entire UAV workflow, from information input to result feedback. Their comparative study reveals a clear gap between academic research, which often emphasizes complex multi-UAV planning and coordination, and industrial practice, which prioritizes practical functionalities such as flight control and human–machine interaction. Feedback from 52 practitioners further highlights concerns about safety, performance, and cost, suggesting that LLMs are currently better suited to supportive and decision-assistance roles rather than direct control in safety-critical UAV systems. In a concise follow-up, Cidjeu et al. [16] reiterate the importance of natural language interfaces, multimodal understanding, and dynamic code generation, and call for a deeper investigation into multi-UAV collaboration and LLM-assisted autonomous decision-making.
Although these surveys primarily focus on UAVs, adjacent domains provide complementary insights. In autonomous driving, Li et al. [62] present a systematic review of the integration of LLMs and MLLMs across perception, prediction, planning, decision-making, and human–machine interaction. Their discussion of multimodal fusion, knowledge distillation, and fine-tuning strategies closely aligns with the challenges encountered in UAV systems. Similarly, Cui et al. [18] review MLLMs for autonomous driving and mapping, summarize datasets and benchmarks, and identify emerging research trends that require coordinated academic and industrial efforts.
Unlike existing surveys, our work offers a focused and in-depth examination of UAV communications, emphasizing how LLMs can enhance communication reliability, autonomy, and decision-making in this critical field. We systematically analyze the LLM functionalities most relevant to UAV communication systems and provide detailed insights into their practical adoption. We also explore the emerging role of MLLMs in integrating visual, sensor, and textual data to enable advanced communication-aware UAV applications. Beyond technical aspects, we address ethical and societal implications, including safety, privacy, accountability, and potential misuse of LAUS. To bridge theory and practice, we present a detailed case study demonstrating how LLM-assisted ICL can support real-world UAV communication scenarios. Finally, we discuss emerging trends and open research directions, positioning our survey as a timely and distinctive contribution to the evolving field of intelligent UAV communications.
Figure 1: Overall organization of the survey, illustrating the research foundations, LLM adaptation stack, core application domains (LLM-enabled communications and MLLM-driven intelligence), and responsible and future intelligence considerations for UAV systems.
I-D Organization of This Survey
The remainder of this paper is systematically structured as follows. Section II details the methodologies for LLM training, adaptation (including fine-tuning, RAG, and prompt engineering), and deployment strategies tailored for UAV operations. Section III explores the application of LLMs specifically in UAV communications, covering navigation, swarm control, and network optimization. Section IV examines the transformative role of MLLMs, detailing their architectures and their application in Vision-Language Navigation (VLN) and human-swarm interaction. Section V critically examines the associated ethical, safety, and societal implications and associated mitigation frameworks. Section VI presents a case study on LLM-assisted ICL for multi-UAV data collection and velocity control. Section VII outlines emerging trends and future research directions, and finally Section VIII concludes the survey. Fig. 1 provides the general organization of the survey.
II LLM Training, Adaptation, and Deployment for UAV Operations
This section presents LLM model architectures and investigates their training, adaptation, and deployment for UAV operations. Training LLMs from scratch enables strong general capabilities, but is elusive for UAV applications due to prohibitive computational, energy, and time costs. Fine-tuning pretrained LLMs is an efficient approach to adapt model architectures to UAV-specific tasks, such as mobility and swarm optimization. However, it requires careful control over overfitting and the use of labeled data. RAG further improves LLM deployment by integrating external knowledge sources with LLM architectures, enabling access to up-to-date operational data and airspace regulations without retraining. Prompt engineering provides a lightweight approach to leveraging existing LLM architectures for rapid deployment and human–UAV interaction, enabling flexible and resource-efficient task optimization.
II-A LLM Model Architectures
II-A1 Encoder-Only Architectures
These are LLM architectures designed to generate rich contextual representations of input data. Their core mechanism is bidirectional self-attention, which enables them to capture a word’s context by considering all surrounding words. This capability makes them particularly effective for tasks that require deep language comprehension, such as understanding objects and scenes. The development of this architecture was pioneered by BERT, which introduced the masked language modeling objective.
In UAV-related applications, encoder-only models have demonstrated strong performance in text understanding tasks. In this line, Silalahi et al. [106] propose a Transformer-based Named Entity Recognition (NER) framework for extracting forensic information from UAV flight logs. Using a manually annotated DJI dataset with six entity types, their BERT-based architecture achieved an F1 score of 91.348%, enabling efficient identification of forensically relevant entities.
II-A2 Encoder-decoder Architectures
These employ a three-module design to handle complex tasks: an encoder that processes the input, a decoder that generates the output, and a cross-attention mechanism that enables the decoder to focus on relevant parts of the input. This structured division of responsibilities often improves the sample and inference efficiency, making it well-suited for models such as T5, BART [58], and OpenBA [132].
The advanced text understanding and generation capabilities of the T5 language model provide a powerful means to enhance the autonomy and utility of UAVs significantly. By enabling natural language command interpretation and automated data summarization, T5 can streamline human-UAV interaction and convert raw sensor data into actionable intelligence, increasing efficiency in critical missions such as disaster response [49]. Encoder-decoder LLMs offer advantages in data efficiency, especially when tasks require both understanding and generative capabilities [71].
II-A3 Decoder-Only Architectures
These are a foundational class of LLM architectures designed for generative tasks, producing human-like text that is coherent and contextually relevant [96]. Their core mechanism is causal (autoregressive) self-attention, which restricts the model to attending only to previous tokens in a sequence when generating the next word. This enables step-by-step text generation while maintaining contextual consistency. The modern era of decoder-only models began with the GPT series, which established the paradigm of pretraining on large-scale text corpora to achieve advanced language-generation capabilities. Today, decoder-only architectures underpin many influential Generative AI (GAI) systems, from research-focused models to leading commercial applications.
In UAV-related applications, decoder-only LLMs have shown substantial potential for natural language interaction and control. For example, Tazir et al. [109] integrated OpenAI’s ChatGPT with the PX4/Gazebo UAV simulation platform, enabling intuitive UAV control through spoken or written commands. Their framework relies on prompt-engineered command generation with a verification mechanism to ensure safety and reliability, achieving high command-generation accuracy in zero-shot settings and demonstrating the feasibility of user-friendly human–UAV interfaces. Additionally, Zhang et al. [134] introduce Open Pretrained Transformers (OPT), a suite of decoder-only models with up to 175 billion parameters. OPT-175B achieves performance comparable to GPT-3 while training with a significantly lower carbon footprint, thereby supporting reproducible research. In general, these studies highlight the versatility of decoder-only LLMs for natural-language-driven UAV control and automation, while also underscoring their reliance on large-scale training data.
Overall, the architecture of an LLM, whether encoder-only, encoder-decoder, or decoder-only, directly affects how effectively it can leverage pre-training to acquire language understanding, reasoning, and generative capabilities. Pretraining converts these architectural potentials into practical intelligence by exposing the model to large-scale general and domain-specific datasets, enabling it to develop emergent abilities essential for UAV applications. The following subsection describes the pretraining process, data preparation strategies, and the challenges of adapting LLMs for UAV-specific contexts.
II-B LLM Pre-training
The pretraining of LLMs primarily relies on next-token prediction, which gives models strong language comprehension, reasoning, and long-range dependency modeling capabilities. In UAV communications, pretraining typically uses two types of data, namely general corpora and UAV-specific datasets. General corpora (e.g., Wikipedia, technical blogs, and web content) provide linguistic diversity and support understanding of technical documentation, whereas specialized datasets focus on UAV- and aerial-specific knowledge. These include $3^{rd}$ Generation Partnership Project (3GPP) standards, Request For Comments (RFC) documents, academic literature, network configuration logs, protocol specifications, simulation reports, and operational UAV repositories. Incorporating such domain-specific sources enhances the model’s understanding of UAV systems, communication architectures, and optimization tasks, making it better suited to UAV-oriented reasoning.
Pretraining LLMs from scratch for UAV applications can be highly resource-intensive. UAV-relevant data, such as flight logs, sensor readings, communication traces, and environmental information, require careful handling in domain-specific formats, including navigation equations, flight dynamics, and swarm coordination rules. Moreover, the computational cost of large-scale training poses significant challenges for real-time UAV systems and edge deployments. For instance, training a frontier-scale model such as LLaMA 3.1 (405B) requires tens of millions of GPU hours and incurs costs on the order of tens of millions of dollars [23]. Consequently, complete pretraining is often infeasible in UAV networks, underscoring the importance of fine-tuning pretrained models to achieve cost-effective and practical domain adaptation.
II-C LLM Fine-tuning
Fine-tuning adapts a pretrained LLM to domain-specific tasks using relatively small datasets, significantly reducing data and computational requirements compared to training from scratch. Using representations learned during large-scale pretraining, fine-tuning enables efficient specialization for UAV networks, improving task performance while maintaining generalization. This makes it a practical approach for deploying LLMs under UAV-specific constraints, such as limited computation and dynamic operating environments.
There are three main fine-tuning paradigms. Unsupervised fine-tuning adapts models using unlabeled domain text, improving domain familiarity but offering limited task precision. Supervised Fine-Tuning (SFT) relies on labeled datasets and often yields strong performance, though annotation costs can be prohibitive in UAV communications. Instruction fine-tuning, commonly implemented through prompt engineering, guides models using natural-language instructions and reduces labeling requirements, although performance is highly dependent on prompt design. Broadly, the fine-tuning plays a critical role in adapting pretrained LLMs to UAV tasks by balancing performance, efficiency, and deployment feasibility [90].
Figure 2: LLM fine-tuning pipeline for UAV applications, showing the progression from a pretrained language model through UAV-specific data preparation, model and training setup, parameter-efficient fine-tuning (e.g., adapters, LoRA/QLoRA, and prompt tuning), followed by evaluation, deployment, continuous monitoring, and updates to produce a UAV-adapted LLM.
As shown in Fig. 2, fine-tuning for UAV applications typically follows a structured pipeline encompassing UAV-specific data preparation, model initialization, training environment configuration, parameter updates, evaluation, deployment, and long-term monitoring. Model parameters may be partially or fully updated, depending on resource constraints, using Parameter-Efficient Fine-tuning (PEFT) [123] to reduce overhead. Here, the PEFT adapts large pretrained models by freezing most parameters and updating only a small subset, such as adapters, low-rank matrices, or prompt embeddings. This strategy substantially reduces training cost, memory footprint, and convergence time while preserving pretrained knowledge and mitigating catastrophic forgetting. Representative PEFT methods include Adapters, Low-Rank Adaptation (LoRA), Quantized Low-Rank Adaptation (QLoRA), Prefix-tuning, Prompt-tuning, and P-tuning, enabling efficient customization of models such as GPT-3, BERT, and LLaMA for UAV applications with limited resources [9].
Overall, fine-tuning offers a practical alternative to full pre-training by reducing computational and temporal costs, though it introduces risks such as overfitting and dependence on labeled data. With techniques such as PEFT, even models with tens to hundreds of billions of parameters can be adapted on modest hardware [107]. In UAV networks, fine-tuning supports task-specific adaptation for applications such as troubleshooting and configuration; however, rapidly evolving domain knowledge often necessitates combining fine-tuning with RAG to maintain up-to-date and context-aware performance.
Table II: Summary of RAG frameworks applied to UAV operation and communication systems, covering decision-making, network optimization, autonomous landing, edge computing, and evaluation methodologies.
Reference Proposed Framework Primary Application Domain Core Technical Approach Key Performance Results
[103] Hybrid LLM–RAG Decision-Making System IoD, Autonomous UAV Missions LLM integrated with RAG over a knowledge graph built from real-time sensor data and historical mission logs; fully auditable decision tracing 92% decision accuracy; 94% mission success rate in dynamic environments
[104] Augmented LLM–RAG IoD Decision Model IoD, Multi-UAV Network Coordination Centralized LLM with RAG processing telemetry, mission logs, and environmental data for real-time contextual decisions BLEU score: 0.82; cosine similarity: 0.87; average decision latency: 120 ms
[11] LLM-Land Autonomous UAV Landing Lightweight LLM with RAG combined with a Vision-Language Encoder (VLE) and Model Predictive Control (MPC) for semantic constraint-aware trajectory replanning High landing success rate; superior obstacle avoidance compared to baseline MPC and non-RAG LLM-MPC; real-time feasibility on constrained hardware
[1] LLM–RAG UAV-BS Optimization Framework UAV-assisted Mobile Networks (UAV-BSs) GAI-assisted LLM with RAG (e.g., GPT-4) to automate mathematical modeling and optimize UAV placement strategies Faster convergence and higher efficiency than Genetic Algorithm-based approaches
[91] GraphRAG General RAG Enhancement (incl. UAV Analytics) Knowledge-graph-based retrieval of nodes, subgraphs, and communities instead of isolated text chunks Improved query-focused summarization; reduced redundancy; enhanced global domain understanding
[100] HybridRAG Domain-Specific Analytics (e.g., UAV Systems) Combination of VectorRAG for semantic similarity and GraphRAG for structured relational reasoning More faithful, comprehensive, and domain-aligned LLM responses
[117] Advanced HybridRAG for UAV MEC Multi-UAV Mobile Edge Computing, Carbon Emission Optimization Integrated KeywordRAG, VectorRAG, and GraphRAG, enabling hybrid factual and relational reasoning Significantly improved reliability and precision in complex optimization model formulation
[133] Enterprise-Grade Modular RAG–LLM Architecture Intelligent Robotic Control Modular RAG pipeline using ChromaDB and Docling with LLMs (e.g., Claude 3.5), grounded in manuals, logs, and safety validators Improved instruction grounding and operational adaptability; challenges in cloud dependence and computational cost
[32] RAGAS RAG Evaluation Reference-free automated evaluation of faithfulness, answer relevance, and context relevance High agreement with human judgment, particularly for faithfulness
[98] RAGCHECKER RAG Evaluation and Diagnostics Fine-grained, component-level evaluation disentangling retrieval and generation quality Stronger correlation with human judgments; effective identification of retrieval–generation trade-offs
Figure 3: Conventional RAG pipeline for UAV applications, illustrating the workflow from UAV task/query formulation through task-aware query encoding, hybrid knowledge retrieval (vector, graph, and hybrid RAG), context augmentation, and grounded LLM-assisted response generation using domain-specific UAV knowledge sources.
II-D Retrieval-Augmented Generation
RAG offers an effective approach to adapting LLMs to the dynamic and rapidly evolving field of UAV networks by grounding generation in external, up-to-date knowledge. Unlike conventional LLMs that depend solely on static training data, RAG retrieves relevant information at inference time, allowing models to incorporate evolving standards, architectures, and transmission technologies critical to UAV networking. Access to authoritative sources such as 3GPP and IEEE standards, emerging network paradigms, and recent advances in signal processing significantly improve the factual accuracy and contextual relevance of the generated outputs.
By grounding responses in curated domain-specific corpora, RAG reduces hallucinations and outdated information without the prohibitive computational cost of training models from scratch or performing extensive fine-tuning. This makes RAG a practical and scalable approach for leveraging general-purpose LLMs in specialized UAV networking tasks [137]. More importantly, evidence-grounded generation enhances trustworthiness by enabling traceable, policy-aligned, and verifiable decision-making. In safety-critical UAV networks, where accuracy, accountability, and compliance are essential, retrieval-assisted grounding underpins the responsible deployment of AI.
A conventional RAG pipeline, as shown in Fig. 3, typically includes several stages: document indexing in vector databases, query processing and embedding, retrieval and ranking of relevant documents, prompt augmentation with retrieved context, and final response generation. These components together balance retrieval accuracy, response latency, and system scalability. Practical deployment also requires alignment with the application context, secure infrastructure, and governance mechanisms to ensure reliable operation.
RAG improves LLM performance in UAV networks by grounding responses in up-to-date, domain-aligned external knowledge. As a result, it offers a cost-effective and flexible alternative to extensive fine-tuning, enabling reliable and adaptive intelligence in rapidly evolving UAV communication and networking environments. For example, Sezgin et al. [103, 104] investigate LLM–RAG–based decision-making frameworks for the Internet of Drones (IoD), demonstrating how grounding LLM reasoning in real-time sensor data and structured mission logs enables context-aware, interpretable, and timely UAV control. Their studies report strong performance in dynamic environments, including up to 92% decision accuracy, 94% mission success rate, and sub-second latency (e.g., 120 ms), highlighting the feasibility of RAG-enhanced LLMs for real-time IoD applications. Beyond high-level decision-making, Cai et al. [11] propose the LLM-Land, a hybrid framework combining RAG-enhanced LLMs with vision–language encoding and model predictive control for autonomous UAV landing in cluttered environments. By translating visual observations into semantic constraints for control, the framework outperforms conventional MPC-assisted baselines while remaining deployable on resource-constrained hardware. Similarly, Afzal et al. [1] applied RAG-enhanced LLMs to UAV-assisted mobile networks, showing that natural-language-driven optimization can accelerate convergence and improve efficiency compared with traditional heuristic methods such as genetic algorithms.
Despite these successes, conventional RAG suffers from inherent limitations, including fragmented retrieval, redundancy, and limited modeling of inter-document relationships. To address these issues, GraphRAG [91] retrieves structured graph elements rather than isolated text chunks, capturing relational dependencies and improving performance on complex reasoning tasks. Building on this idea, HybridRAG combines vector-based semantic retrieval with graph-based relational retrieval to provide richer grounding for domain-specific applications [100]. Wen et al. [117] further extend HybridRAG by integrating KeywordRAG, VectorRAG, and GraphRAG for multi-UAV mobile edge computing, enabling more reliable formulation of complex carbon-emission optimization problems. Related enterprise-grade RAG architectures developed for robotic systems [133] further illustrate how modular retrieval pipelines can reduce hallucinations by grounding LLM outputs in operational logs, manuals, and safety constraints. However, challenges such as cloud dependence, computational overhead, and sensitivity to data quality remain critical considerations when adapting such RAG-enhanced LLM frameworks to UAV systems.
Collectively, these works illustrate a clear evolution: from standard RAG-based UAV optimization, to graph-structured retrieval, to hybrid retrieval systems, and ultimately to fully integrated, enterprise-grade RAG–LLM architectures for autonomous and robotic control. However, evaluating RAG systems remains a significant challenge due to their modular design, the complexity of long-form outputs, and the difficulty of obtaining reliable ground truth. To address this, Es et al. [32] introduced the Retrieval Augmented Generation Assessment (RAGAS). This reference-free evaluation framework measures three core dimensions: faithfulness, answer, and context relevance by focusing on factual grounding, query alignment, and retrieval quality. RAGAS enables rapid, automated evaluation with high agreement with human judgments, particularly in terms of faithfulness. Table II provides a summary of RAG frameworks for UAV operation and communications.
In summary, RAG improves LLM accuracy by grounding responses in external, up-to-date, domain-specific knowledge. However, it increases integration complexity, depends on high-quality knowledge bases, and incurs variable inference costs based on retrieval strategies and database scale. In contrast, prompt engineering offers a lightweight, low-cost, flexible alternative that requires minimal computational overhead and no additional training. Although it requires expertise and is limited by the model’s context window, prompt engineering enables rapid adaptation to diverse tasks and dynamic wireless network environments.
Figure 4: Prompt engineering strategies for UAV applications, comparing in-context learning, chain-of-thought prompting, prompt-based planning, and self-refinement in terms of prompt structure, LLM roles, and resulting capabilities for reliable, scalable, and robust UAV decision-making.
II-E Prompt Engineering: Strategies for UAV Intelligence
Prompt engineering is the process of designing prompts to guide LLMs in generating accurate and useful outputs. As a reusable and adaptable methodology, prompt engineering supports structured problem-solving across diverse domains, improves understanding of LLM behavior, and enables few-shot learning by optimizing in-context prompts. It also plays an important role in the development of advanced UAV systems. As LLM technologies evolve, prompt engineering is expected to become increasingly important, offering finer control over model outputs and enabling new classes of applications.
Creating effective prompts typically involves several key steps. First, the prompt objective must be clearly defined to ensure that the input structure aligns with the intended task. Second, understanding the model’s capabilities and limitations is essential to ensure compatibility between the prompt and the model’s generation abilities. Third, selecting an appropriate prompt format is crucial, as well-structured and explicit prompts consistently lead to higher-quality responses. Additionally, providing sufficient contextual information helps prevent generic outputs and reduces the risk of misinformation, particularly in less-represented or complex scenarios. Finally, prompts should be iteratively tested and refined according to the model’s responses to improve accuracy and reliability before deployment [84].
Prompting techniques are especially valuable for UAV applications that require real-time decision-making and low-latency responses, such as trajectory optimization and resource allocation. These techniques enable LLMs to interpret sensor data and environmental information directly, generating actionable outputs without extensive retraining. This avoids the significant computational overhead associated with conventional model training. For instance, ICL enables LLMs to perform new UAV-related tasks by providing example demonstrations, with performance strongly influenced by the design and organization of these examples. Similarly, CoT prompting promotes deeper reasoning by guiding the model through intermediate steps, enabling it to handle complex multi-stage UAV tasks such as dynamic obstacle avoidance, swarm coordination, and multi-agent decision-making.
Prompt engineering is particularly well-suited to the constraints of UAV networks for three main reasons. First, it is highly resource-efficient, requiring only a forward pass during inference and eliminating the computational and memory costs associated with backpropagation and gradient updates. This makes it suitable for deployment on resource-constrained network devices. Second, it offers high flexibility: by modifying textual prompts, a single LLM can be quickly adapted to a wide range of network tasks without additional coding or retraining. Finally, prompt engineering enables rapid implementation by leveraging the LLM’s inference capabilities to achieve low response times, a critical feature for low-latency services in dynamic UAV network environments.
The core techniques of prompt engineering, including ICL, CoT prompting, prompt-based planning, and self-refinement, are depicted in Fig. 4 and are discussed below.
II-E1 In-Context Learning - ICL
ICL enables LLMs to adapt to target tasks by conditioning on natural-language instructions and demonstrations provided directly in the prompt, without modifying model parameters. In UASNets, ICL operates by providing contextualized demonstrations, such as prior network states, scheduling decisions, and environmental conditions, that allow the model to exploit pretrained knowledge for real-time adaptation. The effectiveness of ICL depends heavily on demonstration selection and formatting, requiring domain expertise in both UAV networking and LLM behavior.
In practical UAV networking scenarios, such as data-collection scheduling, ICL enables systems to dynamically adapt to time-varying network conditions by incorporating real-time feedback into successive prompts. Sensory information, including queue lengths, channel conditions, and battery levels, can be converted into natural-language task descriptions and combined with historical decisions to guide the LLM. This approach is especially useful in emergency scenarios, such as Search and Rescue (SAR) missions, where conventional DRL-assisted methods require extensive training time and computational resources [30].
Emami et al.[30] propose ICLDC, an LLM-assisted ICL framework for emergency data collection scheduling in UASNets. By deploying an LLM at the network edge to generate scheduling decisions from natural-language descriptions of logged sensory data, ICLDC significantly reduces cumulative packet loss relative to DQN and Maximum Channel Gain baselines, while achieving faster convergence and lower computational overhead. To improve robustness, the framework incorporates a safety verifier to override unsafe decisions and an attack-detection module to mitigate jailbreaking attempts.
Building on this work, Emami et al.[29] introduce FRSICL for real-time flight resource allocation in UAV-assisted wildfire monitoring systems. FRSICL jointly optimizes data-collection schedules and UAV velocities using environmental feedback, including AoI, channel conditions, and UAV locations. Simulation results show that FRSICL outperforms PPO and Nearest-Neighbor baselines, achieving lower average AoI and more stable velocity control. Extending to broader disaster-response scenarios, Emami et al.[28] demonstrate that ICL-assisted LLMs can support real-time adaptive decision-making across multiple UAV functions, including path planning, velocity control, scheduling, and power management, while also revealing potential vulnerabilities to prompt-based attacks.
In multi-UAV post-disaster monitoring scenarios, Emami et al.[27] propose AIC-VDS, an attention-assisted ICL framework for joint scheduling and velocity control. By compressing high-dimensional sensory data (e.g., queue lengths, battery levels, and channel conditions) through attention mechanisms, the framework enables an edge-deployed LLM to prioritize critical sensors and generate task descriptions for real-time optimization. Simulation results indicate that AIC-VDS significantly outperforms Multi-Agent DQN and Maximum Channel Gain baselines, achieving up to 91% reductions in packet loss.
Beyond networking-centric tasks, Ji et al.[51] introduce the LLM-assisted Cognitive Reasoning Framework (LLM-CRF) for human–swarm teaming in disaster SAR missions. By integrating ICL with HITL verification, LLM-CRF interprets high-level multimodal commands, decomposes tasks hierarchically, and coordinates heterogeneous UAV swarms. Simulation results show reductions in task completion time, improvements in mission success rates, and lower operator workload, highlighting the effectiveness of ICL for intuitive and safe human–swarm interaction. Table III provides a summary of ICL Frameworks for Emergency UAV and Sensor Network Applications.
Despite these advantages, ICL-based UAV systems remain vulnerable to jailbreaking attacks that manipulate prompts to bypass safety mechanisms. Such attacks can be broadly categorized into white-box methods, which assume access to internal model information, and black-box methods, which rely solely on crafted prompts and are more realistic in deployment scenarios. Addressing these vulnerabilities is essential for deploying ICL in safety-critical UAV networks.
Table III: ICL frameworks leveraging LLMs for emergency UAV and sensor network operations, comparing approaches for data collection, wildfire monitoring, disaster response, multi-UAV coordination, and human–swarm teaming in search and rescue.
Reference Proposed Framework Primary Application Domain Core Function of the LLM Key Advantage over Traditional Methods Key Performance Results
[30] ICLDC UASNets, SAR Emergencies Edge-deployed LLM uses ICL to generate data collection schedules from natural-language task descriptions derived from sensory logs Eliminates iterative DRL training; enables rapid adaptation, improved safety via verifier, and robustness through attack detection Significant reduction in cumulative packet loss vs. DQN and Maximum Channel Gain; faster convergence and lower computational overhead
[29] FRSICL UAV-Assisted Wildfire Monitoring LLM performs real-time flight resource allocation by jointly optimizing data collection schedules and UAV velocities using contextual task descriptions Joint optimization without retraining; improved responsiveness and stability compared to policy-gradient and heuristic methods Lower average AoI and more stable velocity control than PPO and Nearest-Neighbor baselines
[28] LLM-Assisted ICL for Public Safety UAVs Disaster Response and Wildfire Monitoring Edge LLM leverages prompts and demonstrations to adaptively control UAV path planning, velocity, scheduling, and power management Generalizes across tasks without retraining; flexible prompt-based control versus task-specific optimization models Significant packet loss reduction in data collection; highlights resilience and exposure to jailbreaking vulnerabilities
[27] AIC-VDS Multi-UAV Post-Disaster Monitoring (e.g., Tsunami Response) Attention-guided LLM focuses on urgent sensors to generate task descriptions for joint scheduling and velocity control Efficient compression of high-dimensional sensory data; scalable coordination beyond multi-agent DRL limits Up to 91% packet loss reduction compared to Multi-agent DQN and Maximum Channel Gain baselines
[51] LLM-CRF Human–Swarm Teaming for Disaster SAR Central LLM interprets multi-modal human commands, extracts intent, and performs hierarchical task decomposition and mission planning Transforms human role from low-level control to high-level supervision; improves safety with ICL and HITL verification 64.2% reduction in task completion time; 94% mission success rate; 42.9% reduction in operator cognitive workload
II-E2 Chain-of-Thought Reasoning
CoT reasoning represents a significant advance in how AI systems handle complex reasoning tasks by explicitly decomposing problems into intermediate steps [116]. Instead of generating answers directly, CoT prompts models to reason step by step, mirroring human problem-solving processes and improving performance on tasks involving mathematics, logic, planning, and multi-step decision-making. By revealing intermediate reasoning, CoT enhances interpretability and explainability, improves generalization, and helps mitigate hallucinations by constraining generation to verifiable logical sequences.
Since its introduction, CoT prompting has evolved from basic paradigms such as Zero-shot CoT and Few-shot CoT to more advanced and scalable frameworks. These include Auto-CoT, which automatically selects effective demonstrations, and Self-Consistency CoT, which aggregates multiple reasoning paths to improve reliability. More recent extensions, such as Tree-of-Thought (ToT) and Graph-of-Thought (GoT), further structure reasoning as trees or graphs, enabling exploration and comparison of alternative solution trajectories. Although these advanced methods introduce additional computational overhead, they are particularly valuable for complex, intent-driven tasks, such as UAV networking, where transparent, logical, and verifiable decision-making is essential [112].
Zhu et al. [138] demonstrate the effectiveness of CoT in UAV swarm decision-making by integrating an LLM with an RL controller. In their framework, CoT prompting enables the LLM to decompose high-level missions into structured sub-tasks and dynamically adjust the RL reward function based on real-time environmental feedback. A RAG module further enhances situational awareness by fusing multimodal sensor data with knowledge retrieved from prior missions. Simulation results in disaster search and formation navigation scenarios show that this approach outperforms conventional multi-agent RL baselines in success rate, cumulative reward, and sample efficiency, highlighting improved adaptability and coordination in dynamic environments.
Overall, these studies indicate that CoT reasoning enhances adaptability, robustness, and decision-making accuracy in UAV systems operating under uncertainty and rapid dynamics. The progression from simple prompting to structured reasoning architectures positions CoT as a powerful tool for complex UAV tasks where logical transparency and verifiability are critical, even at the cost of increased computational complexity.
However, ICL and CoT prompting are primarily effective for single-task or narrowly scoped scenarios and may struggle with problems that require extensive information integration or long-horizon, multi-stage reasoning. Many UAV networking tasks involve interdependent subtasks, such as coordinated configuration, planning, and execution across multiple agents. Prompt-based planning addresses these limitations by explicitly generating and organizing structured subtask sequences, enabling LLMs to tackle complex objectives more systematically and at scale.
II-E3 Prompt-Based Planning
For large-scale, complex tasks such as network configuration or project development, prompt-based planning breaks down the overall objective into a sequence of manageable sub-tasks. In this approach, the LLM first generates a structured plan of ordered sub-tasks, which are then executed sequentially. This enables the model to address problems that require multi-step reasoning and the integration of information from multiple sources. Prompt-based planning further improves problem-solving by guiding the LLM through explicit action sequences, incorporating intermediate feedback, and continuously refining the plan.
This capability is especially important for UAV operations, where tasks such as formation reconfiguration, resource allocation in multi-UAV systems, and coordinated mission planning require long-horizon reasoning across multiple agents in dynamic environments. Therefore, designing effective prompts and planning structures is essential to the successful deployment of LLMs in real-time UAV control and decision-making applications.
Although prompt-based planning enables LLMs to decompose complex objectives into executable subtasks, the initial plans and decisions produced by the model may still be imperfect or suboptimal. Therefore, self-refinement is essential.
Table IV: Comparison of LLM adaptation techniques—pre-training, fine-tuning, prompt engineering, and RAG—in terms of their core objectives, inputs, processes, resource requirements, advantages, challenges, suitability for UAV applications, and basis of output generation.
Aspect LLM Pre-training LLM Fine-tuning Prompt Engineering RAG
Core Objective Learn general language patterns, knowledge, and reasoning abilities from a massive corpus. Adapt a pretrained model’s knowledge to a specific task or domain using a smaller dataset. Guide a pretrained model’s output for a specific task by carefully designing the input prompt. Enhance a pretrained model’s output by retrieving and incorporating relevant, up-to-date information from an external knowledge base.
Primary Input Massive, general/specialized text corpora (e.g., web data, books, code). Smaller, task-specific, often labeled datasets. A carefully crafted textual prompt (instructions, examples, context). User query + relevant information retrieved from an external knowledge base.
Core Process Unsupervised next-token prediction on a vast dataset. Training from scratch. Updating model parameters (fully or via PEFT) on a targeted dataset. Designing and iteratively refining the input text. No weight updates. Retrieving relevant documents and augmenting the prompt with this context before generation.
Resource Intensity Extremely High. Vast data, immense compute (GPU clusters, 3D parallelism). Moderate to High. Lower than pre-training but still significant for full fine-tuning. Very Low. Requires only a forward pass; no training. Low to Moderate. Adds a retrieval step and a longer context, but inference-only.
Key Advantages Develops foundational language capabilities and emergent skills. High task-specific performance, efficient transfer learning. Highly flexible, fast to implement, resource-efficient. Provides up-to-date & factual information, reduces hallucinations.
Key Challenges Prohibitively expensive/time-consuming. Requires vast domain-specific data. Computationally heavy for edge systems. Requires quality labeled data. Effectiveness depends heavily on the design of prompts. Requires building/maintaining a high-quality, updated knowledge base.
UAV Application Suitability Foundational models via large-scale collaboration (e.g., Federated Learning). Specialized models for mission planning or diagnosis with sufficient data. Ideal for real-time, low-latency tasks (path planning, collision avoidance). Highly promising for dynamic networking tasks with the latest protocols/data.
Output Basis Model’s internal parameters from initial training. pretrained parameters adjusted by fine-tuning data. The model’s inherent capabilities are triggered by the prompt. Synthesis of the model’s knowledge and retrieved external information.
II-E4 Self-Refinement
Self-refinement is an emerging paradigm for improving the initial outputs through an iterative process of self-critique and refinement. In this approach, an LLM evaluates its own outputs and incrementally improves them based on internally generated feedback. For example, if an initial decision, such as a data collection schedule, is suboptimal or incorrect, the same model can be prompted to analyze its own output, identify deficiencies, and generate a revised solution. This significantly reduces the need for human intervention, which is particularly valuable given the complexity of networking and UAV-related tasks [137].
A salient merit of self-refinement is that it requires only a single LLM, which simultaneously serves as the generator, critic, and refiner, without relying on additional training data or RL. The process typically begins with the generation of an initial output, followed by a self-critique stage in which the model provides feedback on its own reasoning or results. This feedback is then used to generate an improved output, and the cycle is repeated until a predefined stopping criterion, such as convergence, task completion, or iteration limits, is met [83].
Building on this concept, Wang et al. [111] propose a closed-loop framework that significantly enhances the reliability and reasoning capabilities of LLMs for autonomous UAV control. Their approach combines a structured prompt framework, Guidelines, Skill APIs, Constraints, and Examples (GSCE), with an iterative feedback and refinement mechanism. After the LLM generates an initial control program using the GSCE framework, the code is executed within a simulator to produce a UAV trajectory. A dedicated LLM-assisted discriminator then evaluates the trajectory against the task objectives and generates targeted diagnostic feedback if deviations are detected. This feedback is subsequently fed back into the GSCE-guided LLM, prompting code refinement and regeneration. The loop of generation, execution, evaluation, and refinement continues until the task is completed or the predefined iteration limit is reached. Experimental results demonstrate that this closed-loop self-refinement system substantially outperforms one-shot generation approaches, achieving task success rates exceeding 90% by enabling the LLM to learn from execution errors and progressively correct its outputs.
A notable merit of self-refinement is its architectural simplicity, as it relies on a single LLM rather than multiple specialized models. However, this simplicity introduces an important consideration: the effectiveness of self-refinement is strongly dependent on the underlying model’s capabilities. Empirical results indicate that performance gains scale with model size and reasoning ability; for instance, when combined with self-refinement, GPT-4 consistently outperforms GPT-3.5 using the same refinement mechanism.
II-F Key Findings and Insights
The adaptation of LLMs to UAV systems requires carefully balancing performance, computational efficiency, and domain specificity. Pre-training offers broad generalization but remains economically and computationally infeasible for most UAV applications. Fine-tuning provides task specialization but depends on high-quality labeled data and introduces risks of overfitting. In contrast, prompt engineering and RAG emerge as lightweight and flexible approaches that enable task adaptation without parameter updates, making them particularly suitable for dynamic UAV environments.
Prompt engineering shapes model behavior through structured instructions and reasoning strategies such as ICL and CoT, enabling real-time decision support with minimal computational overhead. RAG complements this by grounding generation in external, up-to-date knowledge sources, improving factual accuracy in safety-critical UAV operations. While RAG increases system complexity through knowledge base maintenance and retrieval overhead, it reduces hallucination risks and supports continual adaptation without retraining. Importantly, these adaptation strategies should not be viewed as mutually exclusive. Fine-tuning is advantageous for stable, domain-specific behaviors, whereas RAG and prompt-based methods are better suited for rapidly evolving networking standards and mission contexts. Future LLM-assisted UAV systems will likely integrate these techniques within layered architectures, combining lightweight edge models for real-time control with larger models for strategic reasoning and mission planning.
Despite these advantages, effective deployment remains challenging due to constraints such as limited computational and memory resources on edge devices, latency and bandwidth restrictions, long-context processing requirements, reliability and safety concerns, data scarcity, and the dynamic nature of aerial networks. Addressing these challenges necessitates a combination of complementary deployment strategies, including centralized and on-device inference, collaborative and partitioned model architectures, and advances in memory management, long-context handling, chunking mechanisms, and multimodal reasoning.
The following provide a comprehensive discussion of these techniques, examining their trade-offs, recent advancements, and future research directions for scaling LLMs to meet the stringent operational demands of UAVs in mission-critical and highly dynamic environments.
II-F1 Edge, On-Device, and Hybrid LLM Deployment
Thanks to their strong generalization and reasoning capabilities, LLMs enable UAVs to interpret complex and dynamic environments, thereby supporting coherent control strategies for navigation, surveillance, and autonomous decision-making. These capabilities make LLMs particularly attractive for mission-critical, real-time applications such as SAR, Intelligence, Surveillance, and Reconnaissance (ISR), infrastructure inspection, and persistent aerial monitoring. However, most existing LLM deployments rely on cloud-based infrastructure, which has significant limitations in terms of latency, bandwidth consumption, and data privacy. Such constraints render cloud-centric approaches impractical, or even infeasible, for delay-sensitive or communication-constrained UAV operations. Integrating LLMs into UAV platforms has the potential to enhance onboard perception and autonomy significantly, but it also introduces substantial computational and communication challenges. Existing deployment strategies can be broadly categorized into four primary paradigms, as summarized in Table V.
Table V: Comparison of LLM inference paradigms for UAV systems—centralized edge, on-device, split, and collaborative device–server—detailing deployment locations, advantages, challenges, and enabling techniques for real-time, resource-constrained aerial applications.
Inference Paradigm Deployment Location Pros Cons Key Techniques
Centralized Edge Inference Edge server / Ground station Supports real-time decision-making; suitable for safety-critical operations Communication-intensive; relies on stable connectivity Token reduction, model compression, quantization, pruning, parameter-efficient fine-tuning
Local Inference (On-Device) UAV hardware Improves privacy; minimizes latency; robust to connectivity issues Lightweight models lack reasoning depth, high computation, and energy consumption, limited memory/storage Deployment of small models (e.g., OPT-125M/350M), model optimization for embedded platforms
Split Inference (Model Partitioning) Initial layers on UAV, deeper layers on edge/cloud Reduces onboard computation; limits data transmission to intermediate representations Requires coordination between UAV and edge/cloud; latency depends on network Partitioning of LLM layers, sending compact intermediate representations instead of raw sensor data
Collaborative Device–Server Inference Lightweight models on UAV, powerful models on server Reduces bandwidth usage; suitable for unreliable communication; leverages local and remote computation Complexity in coordination; preliminary outputs may require verification Speculative decoding, hybrid device–server collaboration, local preliminary output with remote refinement
  1. a)
    Centralized Edge Inference: In this approach, UAVs transmit sensor data to an edge server or ground station hosting the LLM for inference. Although this paradigm is communication-intensive, techniques such as token reduction can help mitigate transmission overhead. Deploying LLMs at the edge typically requires model compression, quantization, pruning, and parameter-efficient fine-tuning to accommodate constrained hardware resources. When properly optimized, edge-based inference can support real-time decision-making, which is essential for safety-critical UAV operations.
  2. b)
    Local Inference (On-Device): Executing LLMs directly on UAV hardware improves data privacy, minimizes latency, and increases robustness against intermittent connectivity. Prior work, such as Aero-LLM [22], has demonstrated the feasibility of deploying very small models (e.g., OPT-125M/350M) on UAVs. However, such lightweight models often lack the reasoning depth required for complex tasks, including trajectory planning and resource allocation based on ground user requirements. On-device LLM deployment is further constrained by three key factors: (i) high computational demands, where generating even a single token may take several seconds on embedded platforms; (ii) excessive energy consumption, which can rapidly deplete UAV batteries; and (iii) stringent storage and memory limitations, as model parameters and intermediate activations frequently exceed the Random Access Memory (RAM) capacity of typical UAV systems [70].
  3. c)
    Split Inference (Model Partitioning): In split inference, the LLM is partitioned such that the initial layers execute onboard the UAV to process raw sensor data. In contrast, deeper layers run on edge servers or cloud platforms. This strategy substantially reduces onboard computational requirements and limits data transmission to compact intermediate representations, rather than high-resolution raw sensor inputs.
  4. d)
    Collaborative Device–Server Inference: This paradigm leverages cooperation between lightweight onboard models and powerful server-side LLMs. Techniques such as speculative decoding enable a small local model to generate preliminary outputs, which are subsequently verified or refined by a larger remote model. Compared to transmitting raw multimodal data, this approach significantly reduces bandwidth requirements and is particularly well suited to environments with limited or unreliable communication infrastructure [93].
While centralized edge inference can enable real-time LLM processing for UAV applications, its overall effectiveness is fundamentally constrained by the memory capacity, computational resources, and energy availability of edge hardware. These limitations directly influence the scalability and practicality of deploying LLMs in UAV systems. Table V compares LLM inference paradigms for UAVs.
II-F2 Balancing Model Size, Memory, and Performance on the Edge
The scalability of LLMs on edge hardware is fundamentally constrained by memory capacity, computational throughput, and energy availability, all of which are significantly more limited than in cloud environments. A theoretical upper bound on the number of model parameters (P) that can be accommodated on an edge device can be approximated as $P=\frac{M\times 8}{b}$, where $M$ denotes the available memory in bytes and $b$ represents the bit precision per parameter. Under this formulation, an edge device with 8–16 GB of usable RAM can theoretically support approximately 4–8 billion parameters in FP16 precision, or 16–32 billion parameters when employing 4-bit quantization.
In practice, however, this upper bound is substantially reduced by runtime overheads, activation storage, and intermediate buffers, lowering the adequate capacity to roughly 1–10 billion parameters. This practical range is consistent with the scale of current state-of-the-art on-device models, such as Gemini Nano [20] and Qualcomm’s LLMs with up to 10 billion parameters [120]. Consequently, most real-world edge deployments operate in the 1–3 billion-parameter regime, whereas larger models require cloud offloading or distributed execution across multiple devices [72, 36].
Given the stringent memory, computational, and energy constraints of edge platforms, achieving real-time LLM inference on UAVs typically requires not only careful model scaling and optimization, but also high-speed, low-latency communication with edge or cloud servers. 5G networks, particularly through their URLLC capabilities and high-throughput links, provide a critical enabling infrastructure. By enabling rapid computation offloading and seamless access to larger models, 5G connectivity enables UAVs to maintain responsive, context-aware decision-making even in highly dynamic, resource-constrained environments.
Table VI: Techniques for scaling LLMs to extended context windows in UAV applications, comparing selective context filtering, prompt compression, memory mechanisms, and AGI-assisted designs for long-horizon mission planning and interaction.
Technique Purpose Key Mechanism Pros Notes
Selective Context Mitigate long-context processing limitations Filters redundant/low-information tokens using token-level self-information Reduces memory usage and computational overhead; preserves core semantic content; accelerates text generation Minimal performance degradation, but may discard some less obvious informative content
LLMLingua Long-context prompt compression with semantic preservation Hierarchical prompt compression: (i) budget controller assigns compression ratios; (ii) token-level iterative compressor models conditional dependencies Stronger compression than Selective Context; better semantic fidelity; retains critical information More complex; may require additional preprocessing steps; computational overhead during compression
Short-Term Working Memory Maintain local coherence in extended interactions Dynamic buffer for recent observations, actions, intermediate states Ensures short-term continuity; improves immediate reasoning Limited to recent context; cannot store long-term mission knowledge
Long-Term Summary Memory Preserve extended knowledge over long missions Compact, abstracted representation of key patterns, insights, accumulated knowledge Supports strategic reasoning; reduces contextual drift over long durations Abstracted nature may lose fine-grained details
AGI-Assisted LLM Designs Enhance adaptive context retention beyond fixed windows Episodic and semantic memory mechanisms for dynamic storage, retrieval, and updating of information Improves consistency and coherence in long-horizon interactions; adaptive learning Complex architectures; higher computational and design requirements; research-level approach
II-F3 Scaling LLMs for Extended Contexts
LLMs struggle to process long contexts efficiently, which poses a fundamental limitation for real-world applications such as extended UAV missions and long-duration conversational interactions. This challenge stems from the architectural design of Transformer-based models, whose self-attention mechanism scales quadratically with the context length, leading to both computational complexity and memory consumption. Consequently, processing long inputs becomes prohibitively expensive, often necessitating aggressive truncation that can degrade task performance. This limitation is further reinforced during pre-training, as most LLMs are trained on fixed-length context windows, inherently constraining their ability to maintain broader or persistent contextual understanding. Selective Context is an effective strategy to mitigate this issue by filtering out redundant or low-information content before the LLM processes it. By estimating token-level self-information, this approach identifies and discards less informative tokens, thereby reducing memory usage and computational overhead while preserving the core semantic content. Empirical evaluations demonstrate that Selective Context can substantially reduce GPU memory consumption and accelerate text generation with only minimal performance degradation, enabling LLMs to process long documents and extended conversations without severe truncation [67].
A related and more structured approach, LLMLingua, addresses the long-context challenge through hierarchical prompt compression. LLMLingua operates in two stages: first, a budget controller assigns different compression ratios to various components of the prompt (e.g., task instructions, demonstrations) to preserve semantic fidelity; second, a token-level iterative compressor further reduces the prompt by modeling conditional dependencies to retain critical information. Compared to Selective Context, LLMLingua achieves stronger compression while better preserving semantic meaning [52].
Beyond prompt compression, architectural solutions can further alleviate long-context limitations by integrating explicit memory mechanisms. In particular, a conversation memory module can be composed of two complementary components: a) Short-Term Working Memory- A dynamic buffer that maintains recent observations, actions, and intermediate states, ensuring local coherence and continuity; b) Long-Term Summary Memory- A compact and abstracted representation that stores key patterns, strategic insights, and accumulated knowledge over extended mission durations. In addition, Artificial General Intelligence (AGI)-assisted LLM designs can further enhance context retention and adaptive learning capabilities. Unlike conventional LLMs constrained by fixed context windows, AGI-inspired models incorporate episodic and semantic memory mechanisms that dynamically store, retrieve, and update information over time. Such mechanisms reduce contextual drift and improve consistency and coherence in long-horizon interactions and decision-making processes [50]. Table VI compares techniques for scaling LLMs.
II-F4 Structuring Knowledge for Reliable Retrieval
Chunking is a fundamental component of modern RAG systems. As LLM context windows continue to grow, the primary bottleneck has shifted from generation to retrieval, making the structure and granularity of chunks central to overall system performance. Effective chunking breaks information into coherent and meaningful units that reflect how humans naturally organize and understand ideas, rather than relying on arbitrary text boundaries. This distinction is critical because the retrieval quality determines the context the model ultimately sees. Most RAG hallucinations arise from missing, fragmented, or poorly segmented information.
Different chunking strategies, such as fixed-length, hierarchical, semantic, sliding-window, sentence-based, topic-boundary, and hybrid approaches, offer distinct strengths and tradeoffs, and their effectiveness depends heavily on the structure and nature of the underlying content. Ultimately, chunking is not merely a preprocessing step; it functions as an information architecture for AI, shaping how models perceive, retrieve, and reason over knowledge. Better chunking leads directly to more accurate, efficient, and reliable RAG systems.
II-F5 Quantization as an Inference Optimization Technique for Efficient LLMs
Quantization is a widely used model compression technique that improves the efficiency of LLMs by reducing the numerical precision of their parameters, and, in some cases, their activations, from high-precision floating-point formats (e.g., FP32 or FP16) to lower-bit representations such as INT8, INT4, or INT3. This reduction in precision substantially decreases model size, memory footprint, computational complexity, inference latency, and energy consumption, making quantization particularly important for deploying LLMs on resource-constrained edge devices, including Raspberry Pi–class hardware. Quantization can be applied after training via Post-Training Quantization (PTQ), which converts pretrained model weights without retraining, or during training via Quantization-Aware Training (QAT), which simulates low-precision arithmetic to preserve accuracy better. Despite these efficiency gains, quantization introduces tradeoffs: aggressive reductions in bit-width may lead to accuracy degradation, numerical instability, or additional dequantization overhead, particularly for tasks involving complex reasoning or arithmetic. As a result, effective quantization requires carefully balancing efficiency gains with acceptable performance loss to enable practical, sustainable edge AI deployment.
II-F6 Distillation and Prunning
Distillation provides a practical and scalable approach to compressing LLMs by enabling smaller student models to learn from the outputs and internal representations of a powerful teacher, rather than relying solely on task-specific supervision. It can be tightly integrated with techniques such as low-rank adaptation and supernet training, thereby requiring only a small number of additional parameters to train while keeping the pretrained weights frozen. This strategy enables efficient derivation of multiple model variants with different sizes, latencies, and accuracy tradeoffs from a single training run. As a result, distillation not only reduces model size and training cost but also supports dynamic deployment across heterogeneous edge devices, making it particularly well-suited to enterprise and edge AI scenarios where efficiency, flexibility, and performance must be carefully balanced. In parallel, LLM pruning further streamlines models by removing redundant or less important components, producing smaller, faster networks while preserving performance.
In summary, building on the previous discussion of LLM training, adaptation, and architectural choices, it becomes clear that these models provide a flexible and scalable foundation for UAV operations. Due to their high mobility and agile maneuvering, UAVs have been widely adopted for civil and commercial applications, including environmental monitoring, public safety, and parcel delivery. Beyond these roles, UAVs increasingly function as airborne data relays or aerial base stations in infrastructure-limited or disaster-stricken environments, dynamically establishing reliable LoS links with ground sensors to enhance throughput, extend coverage, and reduce operational risk. LLMs can leverage this inherent mobility and the heterogeneous, real-time nature of aerial networks to support communication-aware decision-making, trajectory optimization, swarm coordination, and safety management—tasks that conventional optimization or RL methods often handle inefficiently.
III LLM for UAV Communications
Due to their high mobility and flexible deployment, UAVs have been widely adopted in civil and commercial applications such as environmental monitoring, traffic management, parcel delivery, and precision agriculture [82, 99, 46, 68]. Beyond sensing and service roles, UAVs increasingly function as airborne communication relays for ground sensors in infrastructure-limited or emergency scenarios, including disaster response and border surveillance. By dynamically repositioning to establish short-range line-of-sight (LoS) links, UAVs enable reliable and high-throughput data transmission while extending network coverage beyond fixed terrestrial infrastructure. These capabilities make UAV-assisted networks particularly suitable for remote or hazardous environments, where rapid data collection and wide-area coverage are required with minimal human intervention [31].
LLMs are well suited to UAV communication systems because they directly address key challenges such as dynamic network topology, heterogeneous information, and time-sensitive decision-making. Unlike conventional optimization or reinforcement learning approaches that rely on precise modeling or extensive retraining, LLMs can reason over high-level task descriptions, historical solutions, and real-time feedback to generate communication-aware decisions with low iteration overhead. This section reviews recent efforts to integrate LLMs into UAV navigation, swarm coordination, safety, and network optimization, highlighting their emerging role in intelligent aerial communication systems.
Figure 5: Diagram of the LLM-assisted UAV navigation, planning, and placement framework, illustrating the integration of mission context, reasoning, decision-making, and multi-agent coordination for autonomous operations.
III-A LLM-Assisted UAV Navigation, Planning, and Placement
UAV navigation and mission-level planning are fundamental to autonomous SAR and ISR operations, as they determine where UAVs should move, how coverage should be allocated, and how multiple aerial assets should be positioned to accomplish mission objectives. In complex and dynamic environments, these planning decisions must account for environmental constraints, mission priorities, communication limitations, and regulatory requirements, often under strict time constraints.
Traditional approaches to UAV navigation and multi-UAV planning, such as combinatorial optimization and RL, typically involve high computational costs and long convergence times, limiting their practicality for real-time or large-scale deployment. Other learning-based methods, such as Deep Neural Networks (DNNs) and heuristic algorithms, often lack interpretability and struggle to adapt to rapidly changing mission contexts. Recent advances in LLMs provide a new paradigm for addressing these challenges by enabling high-level reasoning, semantic understanding, and iterative refinement of planning decisions using natural-language task descriptions, historical solutions, and performance feedback. Fig. 5 depicts LLM-Assisted UAV Navigation, Planning, and Placement.
In this subsection, we focus on the use of LLMs at the planning and decision-making layer, including UAV navigation, swarm-level path planning, and multi-UAV placement optimization. Rather than addressing low-level control or execution, the approaches discussed use LLMs to generate, evaluate, and refine high-level plans that guide downstream motion-planning and control modules. This perspective highlights the role of LLMs as cognitive planners that complement traditional optimization and learning techniques in complex UAV missions.
Clader et al.[17] present a novel air–ground robotic system in which a UAV and an Unmanned Ground Vehicle (UGV) collaborate to execute missions specified in natural language within unknown, kilometer-scale environments. Their system employs an LLM-assisted planner, SPINE, to interpret high-level user intent, infer relevant semantic objects, and dynamically generate mission plans. A salient innovation is the construction of a compact, open-set semantic–metric map, built online by the UAV and opportunistically shared with the UGV via the MOCHA communication framework. This design enables effective coordination despite intermittent connectivity. Experimental results in both urban and rural environments demonstrate the system’s ability to complete seven distinct language-specified missions, such as inspecting a blocked bridge or triaging a chemical spill, highlighting its robustness in large-scale, unstructured settings. While this work focuses on air–ground cooperation for single-mission execution, subsequent studies have explored how LLM-assisted reasoning can scale to more complex multi-agent aerial systems.
In this direction, Xiao et al.[121] propose an LLM-assisted framework for UAV swarm path planning in Industry 5.0 scenarios, such as low-altitude rescue missions. The approach addresses the limitations of traditional DNNs and heuristic algorithms in terms of interpretability and adaptability to dynamic environments. The core of the method is the Multi-step Deep Thinking Movement Decision Samples Generation (MSDTMD-SG) algorithm, which creates fine-tuning datasets by simulating human-like multi-step reasoning. Using symbolic modeling and a DRL-inspired reward function, the algorithm identifies optimal movement positions from local observations and embeds the reasoning process into structured templates for LLM training. A cloud–edge collaboration framework enables UAVs to use a cloud-based large LLM when communication is stable and a compressed onboard small LLM when connectivity is limited. Additionally, a scene memory and replay mechanism supports continual learning from failures. Experimental evaluations show that the proposed method achieves the highest task completion rate compared to DRL and Ant Colony Optimization (ACO) baselines, while maintaining competitive computational cost and system latency, and providing superior adaptability in dynamic industrial environments.
Beyond trajectory and path planning, researchers have explored LLM-assisted approaches for high-level multi-UAV decision-making, focusing on placement and coordination rather than low-level swarm control. Wang et al.[114] introduce an LLM-assisted framework for optimizing the placement of multiple UAVs to establish Integrated Access and Backhaul (IAB) networks. Instead of relying on computationally intensive combinatorial optimization or RL methods, the proposed approach employs iterative structured prompting, enabling the LLM to generate and refine candidate UAV positions using problem descriptions, historical solutions, and performance feedback. Simulation results demonstrate that the method achieves near-optimal connectivity and coverage with significantly fewer iterations than conventional techniques, making it suitable for rapid deployment in dynamic environments.
Collectively, these studies demonstrate the versatility of LLMs in enhancing UAV capabilities across a broad spectrum, from single-robot air–ground coordination to swarm-level reasoning and large-scale placement optimization. However, most existing work focuses on isolated components of the autonomy pipeline rather than providing a unified end-to-end solution. This limitation motivates broader efforts to integrate natural-language task interpretation with the full execution stack.
Addressing this gap, Yuan et al.[129] introduce Next-Generation LLM for UAV (NELV), the first end-to-end framework that connects natural-language instructions directly to fully executable UAV flight trajectories across short-, medium-, and long-range missions. NELV integrates five key components, LLM-as-Parser, route planning, path planning, control execution, and real-time monitoring, to enable complex tasks such as multi-UAV patrol, multi-point-of-interest delivery, and long-distance multi-hop relocation. Through three representative use cases, the framework demonstrates how LLMs can significantly reduce pilot workload while adhering to regulatory, airspace, and environmental constraints. Furthermore, the authors propose a five-level autonomy roadmap that outlines the progression from current language-based instruction parsing to fully autonomous LLM-as-Autopilot systems, identifying critical challenges related to regulation-aware reasoning, data scarcity, and safety-critical evaluation.
Recent frameworks such as NELV demonstrate a promising path toward fully autonomous and intelligent UAV operations by connecting natural-language tasking with end-to-end flight execution, including route planning, path generation, control, and real-time monitoring. However, extending these capabilities to large-scale UAV swarms introduces additional coordination and control challenges. Using LLMs for intelligent, scalable, and context-aware swarm coordination offers a compelling approach to addressing these challenges and advancing the next generation of autonomous UAV systems.
Table VII: Comparative overview of recent LLM-assisted UAV planning, coordination, and swarm control frameworks, highlighting their autonomy layers, application domains, core LLM functions, and demonstrated performance benefits across diverse mission scenarios.
Ref. Proposed Framework Autonomy Layer Primary Application Domain Core Function of the LLM Key Results / Advantages
[17] SPINE-based Air–Ground System Mission-Level Planning Air–Ground Cooperative Robotics in Unknown Environments Interprets natural language missions, infers semantic objects, and generates adaptive plans using online semantic–metric mapping Successfully completes language-specified missions in kilometer-scale urban and rural environments under intermittent connectivity
[121] MSDTMD-SG-Based Swarm Planning Planning / Path Optimization UAV Swarm Path Planning for Industry 5.0 Rescue Scenarios Learns human-like multi-step reasoning for movement decisions via fine-tuned datasets and symbolic modeling Highest task completion rate compared to DRL and ACO baselines with competitive latency and improved interpretability
[114] LLM-assisted Multi-UAV Placement Network-Level Planning Multi-UAV IAB Network Deployment Iteratively generates and refines UAV placement using structured prompting and historical feedback Near-optimal connectivity and coverage with significantly fewer iterations than conventional optimization methods
[129] NELV End-to-End Planning Pipeline Natural Language–Driven UAV Mission Execution Maps language instructions to executable flight trajectories via parsing, route planning, control, and monitoring Reduces pilot workload and enables multi-UAV missions while satisfying regulatory and environmental constraints
[81] FlockGPT Execution / Formation Control Natural Language–Driven UAV Swarm Formation Translates high-level geometric commands into executable formation control code using SDFs Enables intuitive swarm control for non-experts; effective formation execution without manual programming
[115] RALLY Execution / Hybrid Control Collaborative UAV Swarm Navigation Combines LLM reasoning with MARL via role assignment and goal refinement Improves task success rate, convergence speed, and robustness over MARL-only baselines
[…]
[13] TypeFly Execution / Low-Latency Control Real-Time LLM-Assisted UAV Control Generates control logic in MiniSpec with stream interpreting for immediate execution Up to 62% reduction in response time and improved real-time responsiveness
III-B LLM-Assisted UAV Swarm Control and Execution
UAV swarms have strong potential for collaborative, distributed missions, but their practical deployment poses substantial challenges at the execution and control layers. In addition to determining high-level plans, swarm systems must translate abstract objectives into real-time control actions, maintain formation and safety, coordinate multiple agents, and respond to environmental uncertainty under strict latency and resource constraints. These challenges are further amplified in scenarios where reliable communication with ground infrastructure is limited or unavailable.
In this context, LLMs have emerged as a powerful interface connecting human intent, high-level reasoning, and low-level swarm execution. By enabling natural language interaction, structured reasoning, and real-time code or policy generation, LLMs offer more intuitive and flexible swarm control compared to traditional interfaces. Unlike planning-oriented approaches, execution-level frameworks emphasize how swarm behaviors are generated, deployed, and coordinated in real-time, accounting for factors such as inference latency, distributed computation, and onboard resource limitations. This subsection focuses on LLM-assisted swarm control and execution mechanisms, including natural language-based formation control, hybrid LLM–RL architectures, distributed LLM inference across multiple UAVs, and low-latency control code generation. These approaches position LLMs not only as planners but also as integral components of the control and system stack, enabling scalable, responsive, and human-centric swarm operation.
Recent research examines how LLMs can improve UAV swarm control by enabling more intuitive and efficient human–swarm interaction. FlockGPT [81] is an early example of natural language–driven swarm formation control. Users provide high-level geometric commands, such as “form a sphere,” which GPT-4 translates into Python code using Signed Distance Functions (SDFs). These SDFs specify target UAV positions, while a flocking algorithm maintains cohesion and prevents collisions. User studies indicate that even non-experts can successfully command and interpret complex three-dimensional formations, demonstrating the accessibility of LLM-assisted swarm interfaces.
Building on this language-driven paradigm, Role-Adaptive LLM-Driven Yoked Navigation for Agentic UAV Swarms (RALLY) [115] introduces a more structured framework that combines LLM reasoning with Multi-Agent RL (MARL). Unlike FlockGPT’s shape-centric control, RALLY focuses on collaborative navigation in dynamic, uncertain environments. It uses a two-stage LLM reasoning pipeline to interpret and refine natural-language goals. Its Role-value Mixing Network (RMIX) dynamically assigns roles – Commander, Coordinator, and Executor – balancing offline LLM knowledge with online MARL policies. This hybrid design improves coordination efficiency, task success rates, convergence speed, and robustness.
While FlockGPT and RALLY focus on high-level planning, SwarmChain [42] addresses a salient systems challenge: LLM inference on resource-constrained UAVs. SwarmChain introduces the CoLLM framework, which distributes LLM computation across multiple UAVs using tensor parallelism and adaptive load scheduling, enabling real-time, on-swarm inference. A control layer employs dynamic prompt templates, vector databases, and context tracking to convert natural-language commands into executable swarm behaviors. Experiments report a 1.9–2.3× reduction in inference latency compared to existing distributed LLM approaches.
In addition, Flying Drones with LLM (TypeFly) [13] addresses latency by reducing the time required for LLMs to generate control code. Instead of using verbose languages such as Python, TypeFly introduces MiniSpec, a compact, token-efficient planning language designed for LLM-assisted control. Features such as skill probes, replanning, and stream interpreting – where actions are executed as code is generated – reduce response time by up to 62%. This significantly improves real-time responsiveness in time-critical UAV operations. Table VII provides a summary of these works.
Overall, these above studies demonstrate the rapid evolution of LLM-assisted UAV swarm intelligence. Each addresses a distinct challenge, including natural language interaction, collaborative reasoning, distributed inference, and low-latency control. Together, they indicate a future where UAV swarms can be commanded and coordinated through scalable, intelligent, and seamless LLM-assisted interfaces.
Table VIII: LLM-enhanced frameworks for UAV safety, network optimization, and autonomous intelligence, comparing methodologies, core LLM roles, and performance improvements in accident investigation, integrated communications, multi-objective optimization, open-world autonomy, and traffic management.
Reference Proposed Framework Primary Application Domain Core Function of the LLM Key Advantage over Traditional Methods Key Performance Results
[125] LLM–HFACS UAV Accident Investigation System UAV Safety Analysis and Accident Investigation LLM reasons over unstructured incident narratives and maps findings to the structured HFACS 8.0 taxonomy to identify causal factors across multiple levels Automates and standardizes human-factor analysis; detects subtle latent errors with reduced time, cost, and human effort compared to manual expert coding Macro-F1 scores of 0.58–0.76 across 18 categories; best F1 of 0.76 (precision 0.71, recall 0.82); some category accuracies >>93%
[75] LLM-assisted Satellite–UAV–IoT Network Optimization 6G Integrated Satellite–UAV–IoT Communications Fine-tuned Llama 3 70B dynamically optimizes device association, resource allocation, and UAV positioning using real-time network telemetry Reduces action-space complexity and decision latency; enables adaptive, predictive network management beyond static optimization and RL methods 27% average spectral efficiency gain; 35% reduction in decision latency compared to traditional approaches
[60] LEDMA (LLM-Assisted Decomposition-Based MOEA) UAV Networks with ISAC LLM acts as a black-box search operator to decompose and solve multi-objective optimization problems for UAV deployment and power control Improves convergence and solution diversity for non-convex, high-dimensional optimization beyond classical evolutionary algorithms Better Pareto front, faster convergence, and higher hypervolume than traditional multi-objective evolutionary algorithms
[136] Embodied Aerial Intelligent Agent Open-World Autonomous UAV Operations Edge-deployed LLM performs high-level task planning and scene understanding, interfaced with low-level perception, mapping, and motion control Enables robust autonomy and generalization in communication-constrained environments through hardware–software co-design 14B-parameter LLM achieves 5–6 tokens/s at 220W; validated across diverse real-world missions (inspection, exploration, monitoring)
[5] LLM-Assisted Unmanned Traffic Management (UTM) System UAV Traffic Management and Airspace Safety LLM analyzes real-time airspace data, checks compliance, and generates actionable alerts and recommendations for human operators Enhances situational awareness and decision support beyond rule-based monitoring systems Up to 91.7% decision accuracy with an average response time of 5.7 seconds across evaluated scenarios
III-C Leveraging LLMs for UAV Safety, Autonomy, and Network Optimization
LLMs are increasingly contributing to various aspects of UAV systems, including safety analysis, autonomous decision-making, communications, traffic management, and education. In safety-critical applications, LLMs can identify contributory factors across multiple causal layers of UAV accidents. In communications, they help optimize the performance of Space–Air–Ground Integrated Networks. LLMs can also be integrated into Unmanned Traffic Management (UTM) systems, where they assist human operators by analyzing real-time data, interpreting complex situations, and generating actionable recommendations. This support improves the reliability, consistency, and timeliness of operational decisions. Additionally, educational platforms that leverage LLMs enable hands-on training in navigation, prompt engineering, and security scenarios, effectively bridging the gap between theoretical knowledge and practical UAV operations.
In the initial research, Yan et al.[125] propose an LLM-assisted UAV accident investigation framework that integrates the Human Factors Analysis and Classification System (HFACS) 8.0 model to automate the analysis of unstructured narrative incident reports. By combining HFACS’s structured taxonomy with LLM reasoning, the system systematically identifies contributory factors across all causal levels, including unsafe acts, latent preconditions, supervisory failures, and organizational influences. The framework was evaluated on 200 UAV incident reports, which contained over 3,600 coded instances, and achieved macro-F1 scores ranging from 0.58 to 0.76 across 18 HFACS categories. The best-performing configuration achieved a macro-F1 score of 0.76, with precision of 0.71 and recall of 0.82, while some category-level accuracies exceeded 93%, matching or surpassing human expert performance on specific tasks. By improving analytical consistency and uncovering subtle latent errors, this approach reduces time and human effort while enabling more comprehensive, data-driven safety improvements.
In the context of communications and network optimization, Liu et al.[75] introduce an LLM-assisted framework for optimizing integrated satellite–UAV–IoT networks in 6G environments. The framework uses a fine-tuned Llama 3 70B model to dynamically manage device association, resource allocation, and UAV positioning based on real-time inputs, including satellite telemetry, IoT device status, and network traffic. To ensure computational efficiency, the authors implement action-space reduction and predictive UAV deployment strategies that adapt to environmental conditions and network load. Experimental results show an average 27% improvement in spectral efficiency and a 35% reduction in decision latency compared to traditional approaches, highlighting the framework’s robustness in dense, dynamic urban scenarios.
Similarly, Li et al.[60] address multi-objective optimization in UAV networks equipped with Integrated Sensing and Communication (ISAC) capabilities. They propose an LLM-assisted decomposition-based multi-objective evolutionary algorithm (LEDMA) that balances communication and sensing performance. By decomposing complex, non-convex optimization problems into simpler subproblems and using LLMs as black-box search operators, the framework iteratively refines solutions for UAV deployment and power control. Numerical evaluations show that LEDMA achieves a superior Pareto front, faster convergence, and higher hypervolume than conventional algorithms, demonstrating the effectiveness of LLMs in solving complex wireless optimization problems.
Beyond optimization and networking, Zhao et al.[136] introduce an aerial intelligent agent that integrates LLM-assisted reasoning with UAV autonomy for open-world task execution. Through a hardware–software co-design, the system deploys a 14B-parameter LLM on an edge-optimized platform, achieving inference speeds of 5–6 tokens per second at a peak power consumption of 220 W. A bidirectional cognitive architecture connects high-level LLM planning with low-level modules for state estimation, mapping, obstacle avoidance, and motion planning. The system is validated across diverse mission-critical applications, including sugarcane monitoring, power grid inspection, mine tunnel exploration, and biological observation. Results demonstrate robust task planning, scene understanding, and generalization, establishing a practical paradigm for embodied aerial intelligence.
LLMs have also been explored as decision-support tools in UTM systems. For example, Atrouz et al.[5] investigate the integration of LLMs into UTM operations, where they assist human operators by monitoring airspace activity, verifying UAV compliance with remote identification and mission plans, and flagging potential violations. Across the evaluated use cases, the system achieved up to 91.7% accuracy and an average response time of 5.7 seconds, demonstrating its potential to enhance safety and efficiency in increasingly congested airspaces.
In addition, the educational domain has also benefited from LLM integration. That is, FLY-LLM Sim [39] is a modular laboratory platform designed for hands-on training with LAUS. The platform includes modules for advanced aerial navigation, where an LLM dynamically selects path-planning algorithms based on natural-language commands, as well as modules for analyzing vision-based cyberattacks on UAV perception systems. By combining the AirSim simulation environment with ChatGPT-4.0 for natural-language command processing and code generation, the platform provides a cost-effective, practical learning experience. User evaluations confirm its effectiveness, highlighting its value for mastering LLM–UAV integration. Table VIII provides a summary of these works.
Figure 6: Architectures for multi-LLM deployment in UAV swarms: parallel/ensemble for robust decision-making, hierarchical for tiered resource-aware intelligence, and distributed specialized agents for decentralized mission execution.
III-D Multi-LLM Architectures for UAV Communications
Single LLM agents are often unreliable for safety-critical systems and can generate biased or unsafe outputs. In UAV swarm operations, where real-time decisions have significant safety implications, relying on a single model increases the risk of errors or unsafe behavior. To address these risks, multiple LLMs can be used to improve reliability, reduce bias, and ensure decisions comply with operational constraints and safety regulations. Multi-LLM systems leverage the complementary strengths of individual models, allowing for specialization, scalability, and robustness. By employing task decomposition and cross-verification, these systems enhance overall decision quality and resilience, thereby making them more suitable for safety-critical UAV swarm applications.
Multi-LLM systems are architectures in which multiple LLMs collaborate, either in parallel or sequentially, to solve complex tasks. Unlike single-model approaches, multi-LLM systems leverage the complementary strengths of individual models, enabling specialization, scalability, robustness, and effective task decomposition. Performance is further improved through cross-verification, which helps detect errors, reduce bias, and mitigate hallucinations. Collaboration among LLMs can follow cooperative, competitive, or ensemble paradigms, while system architectures may be centralized or decentralized, depending on fault tolerance and task allocation requirements. Additionally, multi-LLM frameworks can enhance privacy by routing sensitive queries to trusted models while validating outputs across agents.
Deployment strategies differ significantly between cloud and edge environments. Edge deployments offer low latency and improved privacy by processing data locally, but are constrained by limited computational and energy resources. Therefore, techniques such as model compression, partitioning, and distributed inference are often necessary. In contrast, cloud deployments provide abundant computational capacity and support large-scale reasoning, but may experience higher latency and be less suitable for real-time or privacy-sensitive UAV operations.
In aerial and UAV networks, multi-LLM systems enable distributed coordination among heterogeneous agents, including delivery UAVs, urban air mobility vehicles, and UAV-assisted wireless nodes. Individual LLMs can specialize in tasks such as route planning, communication optimization, or mission-level decision-making, while sharing critical information on navigation, sensing, and system status. This division of labor improves safety, efficiency, and resilience by supporting real-time adaptation, collision avoidance, fault tolerance, and collective optimization – capabilities that are difficult to achieve with a single LLM or a fully centralized controller [80].
Building on this paradigm, Cui et al.[19] propose a UAV swarm inspection system in which multiple lightweight LLM sub-models operate in parallel across a UAV fleet. Local LLMs handle onboard path planning and real-time decision-making, while a centralized orchestrator performs global reasoning, task allocation, and swarm-level optimization. By fusing multimodal data, including maps, telemetry, and visual and thermal imagery, the system dynamically adapts inspection paths and detects defects with high accuracy. Experimental results show improvements in coverage, energy efficiency, and robustness compared to traditional planning methods.
Complementing this approach, Liu et al.[73] introduce a multi-agent LLM framework tailored to UAV-based inspection in the Architecture, Engineering, and Construction (AEC) domain. Their system comprises specialized agents, including Router, PathPlanner, Controller, Perceptioner, and Retriever, that collectively support spatial reasoning, UAV control, and task delegation. A novel image-driven pipeline converts inspection data into semantic point clouds and three-dimensional scene graphs, enabling UAVs to interpret environments in alignment with human intent. Experiments demonstrate improved flexibility and automation in inspection workflows.
Similarly, emphasizing distributed intelligence, Dharmalingam et al. [22] present Aero-LLM, a hierarchical LLM architecture spanning the onboard, edge, and cloud layers. Each layer hosts specialized, fine-tuned models responsible for tasks such as anomaly detection[25], forecasting, and cybersecurity. This hierarchical decomposition enables low-latency responsiveness while supporting high-level analytical reasoning, all within lightweight computational footprints. Evaluations highlight Aero-LLM’s effectiveness in threat detection and secure decision-making on resource-constrained UAV platforms.
Extending multi-LLM coordination to interconnected aerial networks, Yan et al. [126] propose a hierarchical control framework for hybrid terrestrial–non-terrestrial systems incorporating High-Altitude Platform Stations (HAPS). Specifically, a meta-controller LLM deployed on the HAPS manages global resource allocation and load balancing, while onboard LLMs handle real-time motion planning and communication decisions. Experimental results show that this collaborative architecture achieves higher system rewards, lower operational costs, and lower collision rates than DRL-assisted and hybrid baselines. Fig. 6 depicts Multi-LLM architectures for UAV communications, and Table IX provides a summary of the works.
Overall, distributed, cooperative, and domain-specialized LLM architectures represent a promising direction for the future of UAV systems. Multi-agent LLM frameworks enhance reliability, safety, and operational security by enabling robust decision fusion, coordinated resource management, synchronized mission planning, and decentralized control. These capabilities support scalable swarm coordination, real-time decision-making, and adaptive responses in dynamic environments, making multi-LLM systems well-suited for critical applications such as surveillance, disaster response, and infrastructure inspection, where fault tolerance and efficiency are essential.
Table IX: Hierarchical and multi-LLM architectures for coordinated UAV and aerial network operations, comparing distributed, swarm-based, and layered LLM frameworks in inspection, security, and hybrid terrestrial–non-terrestrial network applications.
Reference Proposed Framework Primary Application Domain Core Function of the LLM Key Advantage over Traditional Methods Key Performance Results
[80] Multi-LLM Distributed Coordination Paradigm Heterogeneous Aerial and UAV Networks Multiple specialized LLMs collaboratively handle routing, communication optimization, collision avoidance, and mission-level coordination while sharing critical state information Enables division of labor, real-time adaptation, fault tolerance, and collective optimization beyond single-LLM or centralized controllers Improved safety, efficiency, resilience, and scalability in heterogeneous aerial networks (conceptual and system-level validation)
[19] Multi-LLM UAV Swarm Inspection System UAV Swarm Infrastructure Inspection Lightweight onboard LLMs perform local path planning and real-time decisions, while a centralized LLM orchestrator handles global reasoning and task allocation Balances local responsiveness with global optimization; improves robustness, coverage, and energy efficiency compared to traditional planners Higher inspection coverage, improved energy efficiency, and more robust swarm coordination in dynamic inspection scenarios
[73] LLM-assisted Multi-Agent Inspection Framework AEC Industry UAV Visual Inspection Specialized LLM agents (Router, PathPlanner, Controller, Perceptioner, Retriever) perform 3D reasoning, task delegation, perception, and UAV control Enables spatial-semantic understanding via semantic point clouds and 3D scene graphs; improves automation over rule-based inspection systems Demonstrates flexible task coordination and improved automation of complex inspection workflows
[22] Aero-LLM Secure and Resilient UAV Operations Hierarchical LLMs across onboard, edge, and cloud layers perform anomaly detection, forecasting, and cybersecurity analysis Achieves low-latency response and high-level reasoning while maintaining lightweight onboard computation Strong performance in threat detection and secure decision-making on resource-constrained UAV platforms
[126] Hierarchical LLM-Controlled Hybrid Aerial Network UAV–HAPS Terrestrial–Non-Terrestrial Networks Meta-controller LLM on HAPS manages global resource allocation and load balancing, while onboard LLMs handle real-time motion and communication control Joint optimization of motion and communication outperforms DRL and hybrid control schemes in complex aerial networks Higher system rewards, lower operational cost, and reduced collision rates compared to DRL and hybrid baselines
III-E Key Findings and Insights
III-E1 Integrated Communication–Perception–Reasoning Architectures for Intelligent UAV Communications
Recent UAV platforms integrating 5G connectivity, edge AI, immersive interfaces, and LLM-based reasoning demonstrate a fundamental shift from modular system designs toward tightly coupled communication–perception–reasoning architectures. Rather than treating communication, sensing, and intelligence as independent components, these systems show that their close integration is essential for achieving real-time situational awareness, closed-loop human–UAV interaction, and adaptive decision support in latency-sensitive and bandwidth-intensive environments. This observation indicates that communication must be designed as a first-class enabler of interactive and intelligent aerial systems, rather than a standalone data-delivery service.
III-E2 LLM-assisted Spectrum Management
Recent frameworks for LLM-assisted green spectrum management in 6G Non-Terrestrial Networks (NTNs) integrate LLMs into a hierarchical three-layer architecture comprising data collection, intelligent LLM processing, and feedback-enabled execution. This design enables dynamic and adaptive spectrum allocation across satellites, high-altitude platforms, UAVs, and terrestrial networks. A key innovation is the joint optimization of spectral efficiency and energy consumption, supported by mechanisms including real-time interference analysis, predictive energy modeling, and an explicit energy–spectrum trade-off function. Simulation results show that the LLM-assisted approach significantly outperforms conventional methods, supporting up to 175 active transmitters – a 167% increase in capacity – while maintaining stable energy consumption and tighter interference control under varying threshold conditions. These findings advance intelligent and sustainable spectrum management for future 6G NTNs and underscore the transformative potential of LLMs in enabling green, scalable, high-performance global communications [97].
III-E3 LLM-assisted Near-Field Communications
The integration of LLMs into near-field communications for the Low-Altitude Economy (LAE) brings transformative capabilities that address the unique challenges of UAV-centric wireless networks. By analyzing high-dimensional channel-state information in real time, LLMs can intelligently classify users as near-field or far-field without requiring precise distance estimation, thereby reducing computational overhead. They also enhance multi-user precoding by learning to predict beamforming and power allocation parameters directly from channel data, enabling efficient real-time adaptation to dynamic UAV mobility. Leveraging strong feature-extraction and pattern-recognition capabilities, LLMs can process the massive amounts of antenna and user data generated by extremely large-scale Multiple-Input Multiple-Output (MIMO) systems, thereby enabling performance to scale effectively as network density increases. In addition, their adaptability and few-shot learning abilities enable generalization across diverse operating conditions, such as varying UAV densities or interference levels, without extensive retraining. Finally, LLMs support unified multi-task learning by jointly handling user classification, precoding, and interference management within a single model, significantly simplifying overall system design [124].
III-E4 LLM-assisted ISAC Optimization
The integration of LLMs into the optimization of Integrated Sensing and Communication (ISAC) systems represents a significant methodological advance, as demonstrated by two recent studies that employ LLMs in complementary roles in complex wireless network design. Li et al.[59] examine a multi-UAV ISAC network in which UAVs must jointly provide communication services and radar-based localization of ground users. In this study, an LLM (GPT-3.5 Turbo) is incorporated as an intelligent search operator within a Multi-Objective Evolutionary Algorithm (MOEA). By prompting the LLM with the problem context and current solution states, the model generates novel candidate solutions for UAV deployment and power allocation. This LLM-assisted search enables the evolutionary algorithm to more effectively explore the non-convex trade-off between network utility and sensing accuracy, resulting in improved Pareto fronts and faster convergence compared to conventional approaches.
Complementarily, Li et al.[61] study a multi–base station MIMO ISAC system and address the joint optimization of discrete user association and continuous transmit beamforming. In this framework, the LLM (GPT-o1) serves as a combinatorial solver for the NP-hard user-association subproblem within an alternating-optimization loop. Through carefully engineered prompts incorporating ICL, CoT reasoning, and iterative self-reflection, the LLM infers high-quality association strategies. These results are then combined with rigorous convex optimization techniques—such as fractional programming and the Alternating Direction Method of Multipliers (ADMM), for beamforming design, yielding a hybrid solution that achieves near-optimal performance, closely matches brute-force upper bounds, and outperforms conventional game-theoretic methods.
Overall, these studies highlight a paradigm shift in ISAC optimization, moving beyond purely model-based or data-driven techniques. Instead of serving as standalone solvers, LLMs function as powerful complementary components that enhance traditional optimization frameworks. Their main strength is in addressing problem segments with combinatorial complexity, non-convexity, or a lack of explicit gradient information, where implicit trade-off reasoning is crucial. By encoding domain knowledge and constraints via natural-language prompting, LLMs provide a flexible reasoning interface that reduces reliance on extensive mathematical reformulation and scenario-specific expert tuning.
III-E5 High-Speed Connectivity for LLM-assisted UAVs
The effectiveness of LLM-assisted reasoning in UAV systems depends critically on the reliability, latency, and timeliness of the underlying communication links. While LLMs extend UAV capabilities beyond automation by supporting summarization, semantic understanding, and adaptive mission support, these functions rely on continuous, high-quality data exchange. UAV-assisted URLLC therefore serves as a foundational layer that enables closed-loop perception, real-time human–machine interaction, and dependable reasoning in mission-critical scenarios such as disaster response, search and rescue, and infrastructure inspection. Without ultra-reliable and low-latency communication, the system-level advantages of LLM-assisted intelligence cannot be sustained in real-world deployments.
5G networks, leveraging key technologies such as Millimeter-Wave (mmWave) communication, massive MIMO, and advanced beamforming, enable URLLC and are well-suited for mission-critical applications such as UAV-assisted SAR. URLLC can achieve end-to-end latencies of 1 ms or less, supporting real-time decision-making in highly time-sensitive scenarios. The resulting high-throughput, low-latency communication links between UAVs and edge infrastructure allow rapid transmission of multimodal sensor data for immediate processing. This capability enables on-the-fly LLM inference, allowing UAVs to make responsive, context-aware decisions in dynamic, uncertain environments. By offloading computationally intensive tasks to edge servers, 5G connectivity effectively overcomes many of the resource constraints inherent to onboard UAV platforms.
Nevertheless, despite the advantages of high-speed connectivity, UAV-assisted LLM systems still face intrinsic challenges when handling long data sequences. Transformer-based LLMs have quadratic computational and memory complexity with respect to input sequence length, making long-context processing increasingly expensive even when inference is partially offloaded. As a result, extended missions, continuous perception streams, and prolonged human–UAV interactions remain difficult to support efficiently. Therefore, addressing long-context modeling and efficient attention mechanisms is essential to fully exploit LLMs’ real-time, 5G-assisted capabilities in dynamic UAV operations.
III-E6 Assessing LLM Reasoning and Safety
Field deployments of integrated LLM-enabled UAV systems confirm significant improvements in operational adaptability and effectiveness across diverse applications, including emergency response, environmental monitoring, and infrastructure inspection. At the same time, these evaluations expose persistent constraints related to bandwidth availability, energy consumption, multimodal robustness, and reasoning reliability. These observations highlight the necessity of holistic co-design across communication, computation, and control, as well as the importance of standardized evaluation frameworks, such as UAVBench and UAVThreatBench, to objectively assess reasoning performance and cybersecurity readiness for safe and scalable deployment of intelligent UAV communication systems.
UAVBench marks a significant advancement for the autonomous aerial systems community by providing the first large-scale, standardized, and physically grounded framework for evaluating LLM reasoning in realistic UAV mission scenarios. Prior to UAVBench, no unified methodology existed to assess whether LLMs genuinely understood the critical constraints governing aerial autonomy, such as aerodynamics, flight safety, navigation limits, airspace structure, and ethical decision-making. This absence of standardized evaluation made it difficult to distinguish models capable of supporting trustworthy UAV operations from those that generate superficially plausible but unreliable responses.
The benchmark introduced in [34] addresses this gap by providing 50,000 validated UAV scenarios and an additional 50,000 multiple-choice questions ($UAVBench_{MCQ}$) spanning 10 distinct reasoning styles. This comprehensive design enables systematic evaluation of a wide range of cognitive capabilities, from basic situational awareness to advanced multi-agent coordination and ethics-aware decision-making under operational constraints. Empirical evaluation of 32 state-of-the-art LLMs reveals that, while current models perform relatively well on perception-related and policy-level reasoning tasks, they continue to struggle with ethical judgment, resource-constrained planning, and the complex trade-offs inherent in real-world aerial missions. By releasing both the benchmark and its supporting tools, UAVBench enables the community to diagnose model limitations, compare approaches on a standard foundation, and guide the development of next-generation UAV intelligence through rigorous, transparent evaluation.
Complementing UAVBench, UAVThreatBench introduces a structured benchmark for evaluating LLM performance in UAV cybersecurity within industrial operational technology environments, in accordance with the European Radio Equipment Directive (RED). The benchmark comprises 924 expert-curated scenarios, each annotated with five distinct threats, yielding 4,620 RED-mapped threats across network and device integrity, privacy protection, and fraud prevention. Seven state-of-the-art LLMs, including OpenAI’s GPT-4 variants and LLaMA-3 models, were evaluated on 100 scenarios using fuzzy matching to assess alignment with expert-defined threats. Instruction-tuned models, such as GPT-4o, achieved the highest performance, reaching nearly 87% total threat matching, while open-weight models showed lower alignment and particular difficulty with availability and backend-layer threats. Further semantic analysis revealed systematic gaps in detecting denial-of-service vectors and vulnerabilities in MES/WMS logic.
Together, UAVBench and UAVThreatBench provide complementary, standardized evaluation frameworks that address both operational reasoning and cybersecurity resilience. By offering objective, mission-relevant benchmarks, these contributions equip the UAV research community with tools to assess, compare, and improve LLM-assisted autonomy, paving the way for more reliable, responsible, and deployment-ready intelligent aerial systems.
IV MLLMs: Integrating Vision, Language, and Reasoning in UAV Operations and Communications
This section examines the emergence of MLLMs as a unifying intelligence paradigm that integrates visual, linguistic, and sensor information for UAV systems. By extending large language models with multimodal perception and reasoning capabilities, MLLMs enable UAVs to interpret complex environments, understand high-level human intent, and support context-aware autonomous decision-making.
Recent advances in MLLMs have demonstrated that jointly reasoning over language, vision, and heterogeneous sensor inputs is critical for aerial platforms operating in dynamic, partially observable environments. In UAV applications, this multimodal integration enables intuitive human–machine interaction, robust perception and navigation, and adaptive mission execution by grounding abstract instructions in real-time sensory observations. Compared with text-only models, MLLMs provide a more natural interface between human intent and autonomous aerial behavior, particularly in scenarios involving VLN, swarm coordination, and mission-level planning.
Building on this foundation, this section reviews the core architectural principles of MLLMs, practical training and adaptation strategies for aerial data, and emerging multimodal reasoning mechanisms such as Multimodal Chain-of-Thought (MCoT) and Multimodal In-Context Learning (M-ICL). We then survey representative UAV applications enabled by MLLMs, including human–swarm interaction, VLN, autonomous mission planning, and real-time operational intelligence. Finally, we discuss key challenges related to safety, efficiency, and reliability that must be addressed to enable robust deployment of multimodal intelligence on resource-constrained aerial platforms.
IV-A Building and Fine-Tuning Multimodal LLMs
Training MLLMs for UAV applications begins with constructing multimodal datasets that reflect the unique sensing conditions of aerial platforms. Typical data sources include aerial images and videos, onboard sensor measurements, and associated textual descriptions or operational logs. Compared with ground-level imagery, UAV data exhibit distinctive viewpoints, scale variations, and occlusions, making UAV-specific datasets essential for effective multimodal alignment. Consequently, data augmentation and domain-aware preprocessing are commonly employed to improve robustness under diverse flight conditions. Most MLLMs are initialized through large-scale self-supervised multimodal pretraining to acquire general cross-modal representations. While techniques such as contrastive learning are widely adopted, their primary role in UAV applications is to provide transferable multimodal priors rather than task-specific aerial knowledge. As a result, pretrained models often require substantial adaptation to handle the geometric and semantic characteristics of aerial scenes.
A key challenge lies in aligning heterogeneous modalities for grounded reasoning. Contemporary MLLMs typically rely on modality-specific encoders coupled with a pretrained LLM, enabling language modules to attend selectively to visual features. This flexible architecture serves as a common backbone for UAV perception and reasoning tasks, but its effectiveness in aerial settings depends strongly on downstream adaptation.
Following pretraining, models are fine-tuned on UAV-specific tasks such as aerial image captioning, visual question answering, and object recognition. Dedicated datasets including AeroCaps [8] and HRVQA-VL [38] support fine-grained scene understanding and multi-hop reasoning over UAV imagery. Empirical evidence shows that task-aware fine-tuning is critical: compact MLLMs (e.g., 2B parameters) adapted with UAV data can outperform significantly larger generic models while requiring only a few gigabytes of FP16 memory.
In practice, training large multimodal models from scratch is rarely feasible for UAV systems. Instead, transfer learning and parameter-efficient adaptation techniques are widely adopted. By freezing pretrained backbones and updating lightweight modules such as low-rank adapters, MLLMs can be efficiently aligned with aerial semantics at reduced computational and memory cost, making them suitable for edge-deployed and resource-constrained UAV platforms [38].
Overall, effective MLLM deployment in UAV applications relies less on generic multimodal pretraining strategies and more on UAV-specific data, efficient adaptation, and careful alignment with operational constraints.
IV-B MCoT and M-ICL: Foundations of Multimodal Reasoning
MCoT reasoning extends the traditional CoT paradigm, originally developed for text-based LLMs, to multimodal settings by incorporating diverse data modalities, including images, videos, audio, speech, 3D data, tables, and charts. At its core, MCoT decomposes complex multimodal tasks into a sequence of intermediate reasoning steps, called thoughts, that together form a structured rationale before producing the final answer. Formally, given a multimodal input composed of language tokens $\mathcal{T}$ and non-linguistic modalities $\mathcal{M}$, MCoT can be categorized into two scenarios based on the nature of the rationale $\mathcal{R}$. In Scenario 1, the model generates text-only rationales to reason over multimodal inputs and outputs. In Scenario 2, the rationale itself incorporates multimodal elements, either as purely non-textual signals or as a hybrid of textual and multimodal content. By explicitly modeling the step-by-step progression from perception to cognition, this structured reasoning paradigm enhances interpretability, reduces hallucination, and improves performance on complex multimodal tasks.
The development of MCoT is closely linked to recent advances in multimodal LLMs such as GPT-4V, Gemini, and Claude3, which combine multimodal encoders with autoregressive decoders to process and generate cross-modal content. Despite this progress, several challenges remain, including the computational sustainability of long reasoning chains, error accumulation and propagation across intermediate steps, modality imbalance, hallucination mitigation, and the need for explainable and ethically aligned reasoning frameworks. Future research directions emphasize hybrid neuro-symbolic integration, adaptive reasoning in dynamic environments, and interdisciplinary insights from cognitive science, to advance MCoT toward more general and robust multimodal AGI [113].
M-ICL is another important emergent capability of large multimodal models. By presenting a small number of multimodal examples directly in the prompt, M-ICL enables models to solve novel tasks by analogy, significantly reducing reliance on large labeled datasets. Its training-free nature allows seamless deployment across a wide range of inference-time scenarios. Formally, M-ICL is defined as an inference-time adaptation mechanism for Large Multimodal Models (LMMs), in which the model processes a context $\mathcal{C}$ consisting of $\mathcal{N}$ multimodal demonstration tuples $(I_{i},T_{i},R_{i})$. Here, $I_{i}$ denotes an image, $T_{i}$ represents optional associated text (such as a question or instruction), and $R_{i}$ is the corresponding target response. This context is accompanied by a query $Q=(I_{Q},T_{Q})$, and the model generates an output $O=\text{LMM}(C,Q)$ under a causal language modeling objective, without any updates of gradient-based parameters.
Figure 7: End-to-end multimodal reasoning and control frameworks for VLN in UAVs, featuring instruction grounding, hierarchical collaboration, and safety-aware execution for continuous 3D navigation.
IV-C MLLM-Assisted Human–Swarm Interaction and Autonomy
Integrating MLLMs into UAV systems enables UAVs to more effectively synthesize information from multiple sensors. In the following, we examine how MLLMs can enhance the performance and scalability of UAV systems, focusing on their potential to improve multi-agent perception, navigation, and coordination within UAV swarms, as well as their role in ensuring secure and resilient swarm operations.
Recent advances in MLLMs have significantly expanded the capabilities of Human–Swarm Interaction (HSI) and autonomous UAV systems. Eunmi et al.[33] highlight this shift by introducing SwarmChat, a context-aware interaction framework in which LLM-assisted modules, context generation, intent recognition, task planning, and modality selection collaborate to interpret user commands and adapt them to real-time swarm states. Their results demonstrate that natural-language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, the system’s rule-based intent recognition and limited experimental scope indicate room for more scalable, learning-based solutions. Building on the idea that LLMs can mediate complex swarm behaviors, Liu et al.[77] present a multimodal GPT-4–based approach for UAV formation control using both visual and textual inputs. Their results show that MLLMs can extract commands from human instructions, interpret environmental cues from onboard imagery, and plan formations with over 80% success. This work validates the feasibility of end-to-end multimodal reasoning for swarm perception, coordination, and adaptive task execution, complementing the interaction-focused perspective introduced by SwarmChat. Extending multimodal reasoning into broader autonomous decision-making, Ping et al.[93] explore how MLLMs can enhance multi-agent perception, navigation, and coordination within UAV swarms. Through a forest-firefighting case study, they demonstrate that an MLLM (Doubao 1.5 Vision Pro) can perform natural-language task planning, dynamic fire assessment, and coordinated execution across multiple UAVs. Their discussion of challenges, such as hallucination, inference latency, and communication bottlenecks, further contextualizes the limitations identified in earlier work and underscores the need for more efficient, robust multimodal pipelines.
Table X: Multimodal LLM applications in UAV swarm operations, comparing frameworks for human-swarm interaction, autonomous formation, threat detection, natural language control, and emergency monitoring across diverse sensory and linguistic inputs.
Reference Purpose Input Modalities Models Key Outcomes / Achievements
[33] Human-swarm interaction, context-aware command Text, voice, teleoperation LLM-assisted modules: Context Generator, Intent Recognition, Task Planner, Modality Selector Reduces cognitive load, interprets commands accurately, provides flexible communication modes; requires further learning-based intent understanding and large-scale validation
[77] Autonomous swarm formation, safe command execution Images, text GPT-4 multimodal LLM, pre-training on single-UAV Achieves 82.7% command extraction accuracy; 83.8% formation planning success; handles ambiguous human instructions, integrates perception and swarm status
[93] Dynamic mission planning, multi-agent coordination, situational awareness Text, visual, sensor data MLLM (Doubao 1.5 Vision Pro), chain-of-thought reasoning, multi-source data fusion Autonomous swarm management, adaptive task execution, natural language interaction; addresses hallucination and inference-latency trade-offs
[122] Cyber-physical threat detection in UAV swarms Sensor data, logs, UAV communications Large multimodal models, hierarchical zero-trust architecture, edge-fog-cloud scheduling Improves robustness and timeliness in detecting threats; balances inference accuracy, latency, and computational constraints
[7] Intuitive UAV control via natural language Text, video, image MLLMs (Gemini / LLaVA), logistic regression classifier, YOLOv10, DeepSORT, MAVSDK High-accuracy command execution (97.58% efficiency in video selection), effective autonomous operation, reduces need for specialized programming skills
[127] Real-time fire monitoring, automated report generation UAV imagery, ground sensors, meteorological data LLM (DeepSeek), Grounding DINO, YOLOv8, improved A* path planning Generates combined text-and-image reports, identifies fireground elements, optimizes evacuation routes, enhances situational awareness and operational efficiency
In addition to these operational advances, Xiao et al. [122] examine the role of MLLMs in secure and resilient swarm operations and propose a multimodal intrusion-detection architecture that integrates cyber and physical data streams. Their hierarchical zero-trust framework distributes different model sizes across edge, fog, and command layers, optimizing inference latency and computational cost while enhancing detection robustness. This security-focused perspective underscores the need for trustworthy MLLM integration as swarms become increasingly autonomous and data-intensive. From a control interface standpoint, Balaji et al. [7] introduce PromptPilot, which uses natural language instructions to control UAVs without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA, along with YOLOv10 and DeepSORT for vision-aware tasks, PromptPilot generates executable control code and demonstrates reliable autonomous performance. This work aligns with SwarmChat and Liu et al. in showing that intuitive human–UAV interaction can be achieved through MLLMs, while emphasizing practical deployment and real-world testing. Finally, Yang et al.[127] focus on rapid multimodal situational reporting, integrating UAV imagery, ground sensor data, and meteorological information to generate comprehensive forest-fire intelligence. Their system fuses Grounding DINO, YOLOv8-based models, and the DeepSeek LLM to identify critical environmental features and compute safe evacuation routes via enhanced A. This application illustrates how multimodal perception and language-based reasoning can support high-stakes environmental decision-making, complementing the swarm-control and interaction frameworks previously discussed. Table X provides a summary of these works.
Overall, SwarmChat integrates modules for context generation, intent recognition, task planning, and modality selection, enabling UAV swarms to interpret user commands and adapt to real-time operational states. Experiments show that natural language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, current rule-based intent recognition and limited testing highlight the need for more scalable, learning-based solutions. Building on this approach, PromptPilot enables UAV control through natural language instructions without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA and vision-aware models such as YOLOv10 and DeepSORT, PromptPilot translates user instructions into executable control code and demonstrates reliable autonomous performance. Together, these systems illustrate the potential of MLLMs to mediate complex swarm behaviors, enhance human-machine interaction, and improve the efficiency, scalability, and resilience of UAV operations.
IV-D MLLM-Assisted UAV Navigation and Collaboration
UAV navigation is increasingly framed as a multimodal reasoning problem that requires the tight integration of language understanding, visual perception, and control. Recent advances in VLN reflect this shift, with emerging frameworks enabling UAVs to interpret free-form instructions, reason over complex visual scenes, and execute context-aware flight in previously unseen environments. Collectively, these developments have led to substantial improvements in instruction-following accuracy, generalization, and robustness, accelerating progress toward language-driven UAV autonomy.
Saxena et al. [102] introduce UAV-VLN, an end-to-end framework that reformulates aerial navigation as a question-answering problem. By combining LLM-based commonsense reasoning with open-vocabulary visual perception, UAV-VLN grounds natural-language instructions in real-time visual observations to generate navigation trajectories (Fig. 7). Experimental results demonstrate strong zero-shot generalization across diverse indoor and outdoor environments, highlighting the effectiveness of cross-modal grounding for aerial VLN.
Extending VLN toward safe and reliable execution, Li et al. [66] propose SkyVLN, which integrates LLM-based instruction reasoning with a safety-aware Nonlinear Model Predictive Control (NMPC) module (Fig. 7). By incorporating fine-grained spatial verbalization and path-history memory, SkyVLN improves robustness to ambiguous instructions while enforcing collision avoidance and dynamic constraints during flight. Simulation results show that coupling high-level language reasoning with model-based control enhances both task success and operational safety.
Rather than modifying navigation or control directly, Chen et al. [15] focus on improving language robustness in UAV navigation. Their LLM-assisted Instruction Rephraser (LLMIR) reformulates vague or out-of-distribution user commands into clearer, model-compatible instructions (Fig. 7). This preprocessing strategy significantly improves navigation success under challenging linguistic conditions, offering a practical solution for user-adaptive and reliable human–UAV interaction.
Moving beyond single-agent settings, Wu et al. [118] extend VLN to collaborative scenarios through DuAl-VLN, a dual-UAV navigation framework operating at complementary altitudes. Their AeroDuo system assigns global reasoning and target localization to a high-altitude UAV equipped with a multimodal LLM, while a low-altitude UAV performs fine-grained, collision-free navigation using structured path planning (Fig. 7). Supported by the HaL-13k dataset, AeroDuo significantly outperforms single-UAV baselines, demonstrating the advantages of cooperative reasoning in complex aerial environments.
At a broader level, Liu et al. [76] introduce AerialVLN, a benchmark task that redefines UAV navigation in open, continuous 3D urban spaces. Unlike ground-based VLN, AerialVLN requires UAVs to navigate freely using first-person visual input while handling altitude variation and ambiguous spatial references. This task formalizes key challenges in cross-modal grounding and 3D spatial reasoning that contemporary frameworks such as UAV-VLN, SkyVLN, and AeroDuo aim to address. A comparative summary is provided in Table XI.
Overall, recent VLN advances demonstrate that tightly integrating LLM-assisted reasoning with visual perception, control-aware execution, and multi-UAV collaboration enables UAVs to translate natural-language instructions into real-time, context-aware flight behavior. These developments establish a strong foundation for future research in scalable, language-driven aerial navigation.
Table XI: Comparison of VLN frameworks for UAVs, highlighting innovations in end-to-end reasoning, safety-aware control, instruction robustness, multi-UAV collaboration, and benchmark development for continuous 3D aerial navigation.
Reference Core Innovation / Focus Key Technical Components Key Outcomes / Performance Limitations / Challenges Addressed
[102] End-to-end VLN; reframes navigation as QA; cross-modal grounding for interpretability. Fine-tuned LLM, open-vocabulary visual perception, novel cross-modal grounding mechanism. Significant improvement in instruction-following accuracy & path efficiency; robust zero-shot generalization in diverse indoor/outdoor settings. Minimizes task-specific supervision; outperforms closed-vocabulary baselines.
[66] Integrates VLN with NMPC for safe, precise urban flight. LLM, fine-grained spatial verbalizer, history path memory, NMPC module for control optimization. Substantial improvement in navigation success rate & operational efficiency; ensures safe, constraint-aware flight. Reduces human operator cognitive load and advances explainable autonomous decision-making.
[15] Addresses out-of-distribution (OOD) instructions via an LLM-assisted Instruction Rephraser (LLMIR). Conditional Transformer (VLCT), SentenceBERT, LLM-assisted Instruction Rephraser (LLMIR). 1.39% absolute improvement in task success rate; +1.51% improvement on OOD test set. Enhances reliability with vague/user instructions; improves user-friendliness.
[118] Dual-UAV collaboration at different altitudes for superior navigation. Multimodal LLM (Pilot-LLM), orthographic mapping, multi-stage pathfinding pipeline. Significantly outperforms single-UAV methods; higher success rates & better generalization on unseen environments/objects. Overcomes limitations of single-UAV perspective via collaborative reasoning.
[76] Proposes a new VLN benchmark for UAVs in continuous 3D outdoor urban spaces. Task definition for continuous, open 3D navigation with first-person visual perceptions. Introduces challenges in 3D spatial reasoning, language understanding, and cross-modal alignment. Addresses the gap in aerial agent navigation tasks and focuses on complex urban environments.
Figure 8: Integrated high-level planning and operational intelligence pipeline for UAVs, using multimodal LLMs/VLMs to interpret instructions, infer goals, and generate flight plans, with real-time feedback loops for execution monitoring and situational awareness.
IV-E LLM/VLM-Assisted High-Level Mission Planning, Task Reasoning, and Operational Intelligence
The rapid advancement of multimodal AI is fundamentally transforming UAV capabilities, enabling aerial platforms to perceive complex environments, interpret natural language instructions, and autonomously execute mission-critical tasks. By tightly integrating LLMs with vision-based perception systems, modern UAVs can infer user intent, generate context-aware plans, and adapt dynamically to real-world conditions with minimal human intervention. These capabilities support a wide range of applications, including rapid SAR operations, strategic mission planning, real-time incident detection, and situational analysis. Multimodal pipelines that integrate high-precision object detection, scene-level reasoning, and natural language summarization enable UAVs to convert raw sensor data into actionable intelligence, thereby bridging the gap between human intent and autonomous aerial execution. The overall process is depicted in Fig. 8.
Recent research has increasingly focused on enabling UAVs to autonomously interpret visual scenes, understand natural language, and perform time-critical operations. Yaqoot et al. [128] exemplify this trend with UAV-VLRR, a rapid SAR framework that integrates a multimodal vision–language reasoning system using ChatGPT-4o and Molmo with a high-performance NMPC. By directly translating language commands and aerial imagery into actionable target coordinates and executing point-to-point NMPC control with integrated obstacle avoidance, UAV-VLRR eliminates the need for preplanned trajectories. Experimental results show significant speed improvements over both human pilots and conventional autopilots, highlighting the effectiveness of combining MLLMs with real-time control for agile, safety-critical missions.
Complementing rapid execution at the control level, Sautenkov et al. [101] extend multimodal reasoning to strategic mission planning with UAV-VLA. Instead of directly controlling low-level flight, their system enables users to specify large-scale aerial missions using natural language, which an LLM interprets as high-level objectives. A vision–language model then identifies relevant targets in satellite imagery, and a final LLM module generates a complete flight plan. Although the resulting trajectories are slightly longer than those designed by expert operators, they are produced more than six times faster, demonstrating the potential of MLLMs to reduce operator workload and streamline global-scale mission design significantly. Beyond navigation and planning, multimodal AI has also demonstrated strong potential for real-time incident detection and situational awareness. Eesaar et al. [24] present an autonomous UAV system that integrates a DJI Matrice 30T with an end-to-end multimodal pipeline for emergency response. In this system, YOLOv12n detects accidents and fires in real-time aerial video, LLaVA-OneVision-Qwen2 generates detailed scene-level descriptions, and GPT-4o mini produces concise natural-language summaries for first responders. Field validation shows that this pipeline can rapidly transform raw aerial imagery into actionable alerts, significantly improving response times and situational awareness in transportation and emergency management scenarios.
Taken together, these works illustrate a growing ecosystem of multimodal UAV systems that translate vision and language understanding into meaningful action. From fast, reactive control in UAV-VLRR, to large-scale mission planning in UAV-VLA, to real-time scene interpretation for safety-critical response, recent advances highlight the expanding role of MLLMs in enabling intelligent, adaptive, and resilient aerial autonomy. As multimodal AI matures, it is set to play a central role in developing reliable, interpretable, and fully autonomous UAV operations across diverse real-world applications.
IV-F LLM/VLM-Driven Autonomous Control and Real-Time Flight Decision-Making
LLMs and VLMs in UAV systems pose a critical challenge: enabling autonomous agents to accurately infer human intent, interpret complex, dynamic environments, and make safe, context-aware decisions despite the uncertainty and potential hallucinations inherent in modern foundation models. Recent advances address these challenges by combining structured prompt execution mechanisms with lightweight, task-specific vision models, improving reliability and reducing erroneous reasoning. Multimodal architectures further enhance human–machine collaboration by generating navigation commands with natural-language explanations and confidence estimates, enabling operators to supervise and guide UAV behavior in real time. Meanwhile, efficient onboard autonomy frameworks leverage compact, fine-tuned LLMs and advanced perception modules for object detection and depth estimation, enabling robust autonomous decision-making even under stringent computational constraints.
As multimodal large models continue to expand the capabilities of autonomous UAVs, recent research has increasingly focused on improving reliability, transparency, and resource efficiency in real-world deployment. Hu et al. [45] address one of the most pressing challenges—hallucination in LLMs and VLMs—through LLVM-UAV, a framework that tightly couples a Domain-Guided Structured Prompt Execution Framework (DGSPEF) with lightweight, task-specific vision models. By translating natural-language instructions into structured, domain-constrained executable code and validating perception using independent visual modules, LLVM-UAV enables reliable zero-shot task execution without additional model training. Extensive hardware experiments and evaluations across eight LLMs demonstrate strong robustness, positioning the framework as a practical solution for safety-critical UAV applications such as mapping, disaster response, and precision agriculture.
Building on the theme of reliability while emphasizing human collaboration, Krupavs et al. [55] introduce a multimodal architecture for human–machine collaborative UAV navigation. Their system uses GPT-4.1-nano for multimodal reasoning and generates navigation commands along with natural-language explanations and confidence scores, which are presented to operators through a touch-based approval interface. This design enables a closed-loop interaction paradigm in which operators can accept, reject, or modify model-generated actions. Experimental results show high levels of user trust and explainability, with approximately 90% approval of AI-generated instructions. However, observed response latency and imperfect calibration between confidence scores and actual performance highlight ongoing challenges in deploying explainable VLMs for real-time aerial control.
While LLVM-UAV and human–machine collaboration frameworks prioritize reliability and transparency, Wang et al. [110] focus on enabling efficient onboard autonomy for resource-constrained UAV platforms. Their LMUCS framework integrates a compact, LoRA-tuned LLM for natural language parsing with advanced perception modules, including YOLOv11 for object detection and MiDaS for monocular depth estimation. Achieving over 94% command-parsing accuracy and strong visual perception performance, LMUCS enables lightweight UAVs to execute complex missions, such as autonomous material search and delivery, without relying on cloud-based computation. This work demonstrates that high-level language reasoning and robust 3D perception can be jointly realized on embedded platforms.
Extending these ideas to goal-directed navigation, Zhang et al. [135] propose VLFly, a unified vision–language navigation framework explicitly designed for UAVs that avoids the common reliance on discrete action spaces and external localization systems. Instead, VLFly directly outputs continuous velocity commands from monocular camera input, integrating an LLM-assisted instruction encoder, a VLM-based goal retriever, and a trajectory planner for real-time waypoint generation. Evaluated across diverse simulation environments without task-specific fine-tuning, VLFly consistently outperforms previous approaches. Real-world flight experiments in both indoor and outdoor settings further confirm its robust open-vocabulary grounding and its ability to interpret abstract or indirect language instructions, highlighting its suitability for zero-shot deployment in unconstrained UAV navigation scenarios. Table XII summarizes challenges and solution directions in LLM/VLM-Based UAV Autonomy.
Integrating LLMs and VLMs into UAV platforms requires addressing challenges in intent understanding, environmental perception, and safe decision-making under model uncertainty. Recent approaches address these issues through structured prompt execution, lightweight task-specific perception, and explainable multimodal interaction. Efficient onboard autonomy frameworks demonstrate that compact language models and advanced visual perception can enable robust, real-time operation on resource-limited UAVs. These advances represent a significant step toward reliable, interpretable, and fully autonomous multimodal aerial systems.
Table XII: Key challenges in LLM/VLM-based UAV autonomy—including model uncertainty, interpretability, resource constraints, environmental perception, and intent understanding—paired with solution approaches, required capabilities, and desired operational outcomes.
Core Challenge Solution Approach Key Capability Desired Outcome
Model uncertainty and hallucinations of foundation models Combining structured language reasoning with robust visual perception Error-resilient, zero-shot task execution across diverse missions Greater reliability and safety in real-world aerial missions
Lack of interpretability and trust in AI decisions Pairing VLM-based reasoning with real-time depth perception and operator-guided approval loops Enhanced explainability and human-AI collaboration Improved transparency and user-aligned operation
Limited onboard computational resources for AI models Using compact, finely tuned LLMs with efficient visual detection and depth estimation pipelines High-level decision-making on resource-constrained UAVs Widely deployable AI-powered UAV platforms
Complex environment perception and context-aware decision-making Integrating domain-guided prompt execution with lightweight vision modules Safe, context-aware decisions in dynamic environments Intelligent operation in complex scenarios
Human intent understanding and mission execution Systems combining language reasoning with visual perception and human-centric interaction Understanding human intent and complex environments More intelligent, interpretable, and reliable systems
Figure 9: Multimodal LLM/VLM framework for semantic understanding and analysis in UAV applications, integrating diverse sensor data for scene reasoning, threat assessment, and domain-specific insights.
IV-G Semantic Understanding, Analysis, and Interpretation
As UAVs increasingly operate in complex environments and regulated airspaces, traditional perception pipelines struggle to deliver the semantic understanding and real-time reasoning required for robust autonomy. Emerging multimodal foundation models offer a promising alternative, enabling UAVs to perform advanced inference, threat assessment, and scene understanding across diverse tasks. Applications include low-altitude airspace security, where data from multiple sensor modalities inform defensive decision-making, and precision agriculture, where vision–language models convert aerial imagery into actionable crop intelligence. These advances highlight the growing versatility of multimodal AI in UAV systems, supporting adaptive, efficient, and generalizable environmental intelligence. Fig. 9 visualizes LLM/VLM-assisted semantic Understanding, Analysis, and Interpretation.
As multimodal large models gain traction in UAV-related domains, recent research has begun to explore their potential beyond navigation and control, extending into higher-level reasoning for security and remote sensing applications. Lei et al. [57] exemplify this shift in low-altitude airspace security by proposing a system that integrates MLLMs with heterogeneous perception data to infer the intent of non-cooperative UAVs. By fusing inputs from RGB cameras, infrared sensors, and radar, the system constructs structured representations, which are processed by an MLLM that leverages contextual knowledge and generative reasoning to assess adversarial behavior and recommend defensive actions. A confrontation scenario demonstrates the feasibility of this approach. The authors also identify key challenges – including robust communication, scalability to multi-UAV scenarios, and effective edge–cloud collaboration – that outline future directions for intelligent and adaptive airspace defense architectures.
Extending multimodal models to environmental monitoring, Bie et al. [10] investigate the use of VLMs for crop segmentation in UAV-assisted remote sensing. Their approach reframes segmentation as a text-driven coordinate prediction task by converting ground-truth masks into polygon-based XML representations. This formulation enables VLMs, fine-tuned with parameter-efficient LoRA techniques, to generate segmentation outputs directly through language. Experimental results show that Qwen2-VL-7B achieves the best performance among the evaluated models, particularly for tobacco segmentation, yielding competitive F1 scores compared with specialized vision-based methods. Although overall accuracy remains lower than that of task-specific architectures, the proposed method offers significant advantages in flexibility and generalization, positioning VLMs as promising tools for scalable and versatile agricultural remote sensing pipelines.
Overall, multimodal AI is rapidly expanding UAV capabilities in both security and environmental monitoring. By integrating LLMs with visual and sensor data, UAVs can interpret complex scenes, infer intent, and make context-aware decisions in low-altitude airspace. At the same time, VLMs facilitate flexible environmental analysis, such as text-driven crop segmentation for precision agriculture. Together, these developments highlight the potential of multimodal AI to enable intelligent, adaptive, and generalizable UAV operations.
IV-H Key Findings and Insights
This section consolidates key observations and empirical findings on the deployment of Multimodal Large Language Models (MLLMs) in UAV applications. Emphasis is placed on quantitative evidence, benchmark results, and UAV-specific limitations that directly affect real-world reliability, efficiency, and safety as sequel.
IV-H1 Benchmarking MLLMs for Multi-UAV Collaborative Perception
AirCopBench [131] is the first benchmark explicitly designed to evaluate MLLMs in multi-UAV collaborative perception under degraded sensing conditions. The benchmark contains over 2,900 multi-view aerial images and more than 14,600 VQA samples, covering four task categories (Scene Understanding, Object Understanding, Perception Assessment, and Collaborative Decision) across 14 task types. Evaluation of 40 state-of-the-art MLLMs reveals a substantial performance gap: the best-performing model achieves only 59.17% accuracy, lagging human performance by more than 24% on average, with large variance across task categories. Fine-tuning results indicate that sim-to-real transfer can partially mitigate these gaps, highlighting the importance of UAV-specific data and training strategies for collaborative aerial perception.
IV-H2 Speculative Decoding for Efficient LLM Inference
Inference latency remains a critical bottleneck for UAV deployment, as conventional autoregressive decoding requires one forward pass per generated token. Speculative decoding improves efficiency by allowing multiple candidate tokens to be generated and verified in parallel within a single forward pass, significantly reducing end-to-end latency while preserving output correctness. Recent approaches such as EAGLE-3 further eliminate the need for a separate draft model by exploiting internal hidden states of the target model. Although primarily evaluated in general-purpose LLM settings, these techniques are highly relevant to UAV platforms, where real-time responsiveness and energy efficiency are tightly constrained.
IV-H3 Long-Context Modeling in Multimodal Systems
Long-context modeling enables multimodal systems to reason over extended inputs, but scaling context length introduces challenges including quadratic attention complexity, positional encoding degradation, and mid-context retrieval failure. Recent solutions—such as efficient attention mechanisms, memory-augmented architectures, and chunked processing—improve scalability, while data-centric strategies (e.g., synthetic long-context generation and short-to-long alignment) enhance robustness. However, for UAV applications involving long-duration missions or multi-stage reasoning, a gap persists between theoretical context length and effective usable context, limiting practical long-horizon multimodal reasoning [78].
IV-H4 Multi-Agent Framework for MLLMs Training
The high computational cost of end-to-end multimodal training motivates modular and multi-agent alternatives. BEMYEYES [47] adopts a collaborative framework in which a lightweight visual “perceiver” extracts scene information and a powerful text-only LLM acts as a “reasoner.” Through multi-turn interaction, this system enables multimodal reasoning without training large-scale vision–language models. Experimental results show that this fully open-source framework can match or surpass proprietary models such as GPT-4o on several knowledge-intensive tasks, demonstrating a scalable and resource-efficient path for multimodal reasoning in UAV systems.
IV-H5 Addressing Visual Hallucination in Multimodal LLMs
MLLMs have demonstrated remarkable capabilities across a wide range of tasks that integrate text and visual information. Despite these advances, a persistent challenge is Visual Hallucination (VH), the generation of outputs containing details inconsistent with the visual input. In other words, MLLMs may ”imagine” objects, attributes, or relationships that are absent from an image, particularly in tasks such as object recognition. This phenomenon undermines model reliability and limits their safe and effective deployment in UAV applications.
IV-H6 Balancing Efficiency and Reasoning Diversity in Vision-Language Models for UAV Applications
Results in [38] demonstrate that UAV-oriented vision–language models can significantly outperform general-purpose models when adapted with domain-specific data and training strategies. The proposed UAV-VL-R1 model achieves strong reasoning performance while maintaining a compact footprint suitable for onboard deployment. Ablation studies reveal a key trade-off: while supervised fine-tuning (SFT) accelerates convergence and stabilizes early performance, excessive reliance on SFT reduces reasoning diversity, particularly in counting and numerical inference tasks. In contrast, the multi-stage GRPO framework enhances reasoning diversity and cross-task generalization through intra-group reward comparisons. The associated HRVQA-VL dataset, consisting of 50,019 samples organized into progressively challenging stages, further highlights that domain-specific data can outweigh sheer model scale, outperforming both 2B and 72B Qwen2-VL baselines under UAV constraints.
IV-H7 Understanding and Evaluating Hallucination Behavior in Aerial LVLMs
Basak et al. [8] introduce AeroCaps, the first aerial-view image captioning dataset, comprising 1,256 human-annotated images. Their analysis shows that aerial perspectives significantly amplify hallucination rates, especially for object identity, spatial relations, and numerical descriptions, with error frequency increasing for longer captions. Notably, mitigation strategies such as object hints and LLM-based self-evaluation exhibit limited effectiveness, underscoring the difficulty of reliable hallucination detection. These findings indicate that current LVLMs remain insufficiently robust for real-world UAV deployment without further advances in dataset scale, evaluation rigor, and grounding-aware model design. The study emphasizes that, despite impressive performance on conventional benchmarks, current LVLMs are not yet sufficiently robust or trustworthy for deployment in real-world aerial applications, including automated surveillance, environmental monitoring, and disaster response. The results suggest that progress in this area will require not only improved model architectures and training strategies but also domain-specific datasets, rigorous evaluation protocols, and transparent reporting of failure modes. As aerial data becomes increasingly central to AI-driven decision-making, addressing these challenges will be critical to ensuring the safe deployment of AI.
In summary, based on Sections II and III, the multifaceted integration of LLMs and MLLMs into autonomous aerial systems necessitates an investigation of their ethical implications. LLMs are becoming central to UAV decision-making, perception, and control – capabilities with significant real-world consequences. Section II describes how LLMs, through prompt engineering and RAG, enable UAVs to perform complex tasks such as disaster response and navigation autonomously. The shift to MLLMs in Section III further deepens this integration, allowing UAVs to fuse and interpret streams of visual, sensory, and linguistic data to navigate urban environments, manage swarms, and conduct mission-critical operations such as SAR. Additionally, the technical capabilities and inherent vulnerabilities of LLMs, as outlined in these sections, create specific and urgent ethical imperatives. For example, LLMs are prone to hallucinations, generating plausible yet nonfactual content, which raises concerns about their reliability. This flaw could lead to significant navigation errors or misidentification of targets in the field. Their black-box nature and sophisticated reasoning make it difficult to audit critical UAV decisions. This opacity is compounded in MLLMs, where adversarial attacks can exploit visual vulnerabilities to bypass safety protocols and potentially hijack UAV systems. Moreover, LLMs are trained on large-scale datasets that inevitably contain societal biases; if a model prioritizes disaster relief efforts or the identification of persons of interest, it could systematically disadvantage certain groups. Finally, there are environmental concerns regarding energy consumption and sustainability. Therefore, investigating the ethical implications is critical to the responsible development and deployment of LLM-assisted UAVs. This ensures that the pursuit of technological autonomy is guided by safeguards for safety, fairness, accountability, and societal trust, preventing harm and aligning these systems with human values and safety-critical standards.
V Ethical Challenges and Best Practices for LLM Integration in UAVs
As LLMs and MLLMs become integral to autonomous UAV decision-making, their deployment introduces critical ethical and safety challenges that must be systematically addressed. Privacy risks arise when large-scale datasets include sensitive information collected without explicit consent. Bias and fairness issues may emerge from imbalanced training data, leading to unsafe or discriminatory UAV decisions. Accountability and transparency are further complicated by the black-box nature of LLMs, which hinders failure diagnosis and responsibility attribution. In addition, the high computational cost of LLM training and inference raises sustainability concerns, motivating energy-efficient and environmentally responsible AI practices.
This section reviews representative ethical challenges in LAUS and surveys emerging best practices for secure, transparent, and accountable deployment in real-world UAV scenarios.
V-A Securing UAV Operations: Privacy and Data Protection with LLMs
While LLMs enhance autonomy and reasoning in UAV networks, they also introduce security and privacy vulnerabilities, including prompt manipulation, leakage of sensitive sensor data, unauthorized access, and unsafe decision-making under adversarial conditions. When LLMs influence mission planning, resource scheduling, or control actions, such vulnerabilities can result in severe operational failures. LLM security encompasses practices that protect models, data, and outputs across development, deployment, and real-time operation. Core objectives include preventing unauthorized access, mitigating adversarial attacks such as jailbreaking and data poisoning, and ensuring compliance with safety and operational constraints.
Jailbreaking is a particularly critical threat in LAUS. These attacks manipulate prompts to bypass safety constraints and induce unauthorized behavior. White-box attacks assume access to internal model information, enabling gradient- or logit-based manipulation, while black-box attacks rely solely on crafted prompts and are therefore more practical in real-world deployments. Among black-box methods, template completion attacks are especially dangerous. By embedding harmful intent within structured templates, fictional scenarios, or role-playing contexts, attackers exploit LLMs’ pattern-completion behavior while evading keyword-based filters. Context-based manipulation and code injection further amplify these risks, particularly for models relying heavily on contextual coherence and demonstration-based reasoning. These risks are amplified in UAV networks using ICL for real-time decision-making. For example, malicious demonstrations injected into scheduling prompts can bias UAVs toward suboptimal sensing actions, such as prioritizing sensors with poor channel conditions, leading to increased packet loss, degraded network performance, and compromised mission objectives [30]. This illustrates how prompt-level attacks can propagate into system-level failures.
Mitigation requires a multi-layered defense strategy. Integrating an independent safety verifier allows LLM-generated actions to be evaluated against predefined constraints and overridden if unsafe. Complementary attack detection modules can monitor prompt statistics, such as perplexity or structural anomalies, to identify manipulations in real time. Together, these mechanisms enable robust operation under adversarial conditions.
Table XIII: Framework for LLM observability and safety verification in UAV decision-making systems, detailing monitoring mechanisms, traceability, feedback integration, and rule-based verification to ensure transparency, reliability, and operational safety.
Concept Goal Key Mechanism Benefits Notes
LLM Observability Ensure trustworthy, explainable, and consistent outputs Continuous monitoring of model outputs using defined evaluation metrics (correctness, relevance, coherence, factual consistency) Detects model drift, data shifts, hallucinations; enables optimization and transparency Requires systematic logging, monitoring infrastructure; basic metrics alone are insufficient
Prompts and User Feedback Track quality and relevance of LLM outputs Structured logging of prompts, API calls, and user feedback annotated with metadata Identifies prompt failures; optimizes prompt templates; improves user experience Dependence on accurate user feedback; may require additional preprocessing
Tracing Understanding system performance and workflow Captures traces of user interactions, broken down into spans representing workflow steps Pinpoints bottlenecks; shows how components interact; supports debugging Requires proper instrumentation; may add computational overhead
Latency and Usage Monitoring Optimize performance and resource utilization Measure response times, token usage, throughput, and resource consumption Identifies bottlenecks; helps scale efficiently; reduces unnecessary latency Does not capture quality or correctness of outputs by itself
Safety Verification Layer Ensure reliable and safe decision-making Independent verifier applies rule-based scoring against predefined safety constraints; overrides unsafe LLM outputs Guarantees safety compliance; prevents execution of unsafe or illogical actions Introduces additional processing; relies on well-defined safety rules; deterministic rules may limit flexibility
LLM + Safety Verifier Integration Combine reasoning flexibility with safety LLM generates adaptive decisions; safety verifier enforces constraints before execution Balances adaptability and reliability; ensures safe operation in dynamic systems Coordination between LLM and verifier needed; probabilistic LLM outputs still require careful monitoring
V-B Bias-Aware LLMs for Safer UAV Operations and Communications
Bias in LLMs arises from imbalanced or unrepresentative training data and can directly affect safety and mission effectiveness in UAV systems, for example through skewed threat assessment or unfair scheduling decisions. These risks are particularly severe in post-disaster scenarios characterized by uncertainty, limited communication, and strict latency constraints. In large-scale SAR missions, relying on a single monolithic LLM for perception, planning, and scheduling can create an operational bottleneck. Under high cognitive load, such models are more prone to hallucinations, biased prioritization, and delayed responses, violating key requirements for multi-UAV operations, including real-time responsiveness, efficient resource use, and robustness to partial failures.
To address these limitations, recent LLM-assisted UAV architectures distribute reasoning across ensembles of specialized agents responsible for distinct tasks, such as trajectory planning, sensing scheduling, or risk assessment. Bias can be further mitigated by supervisory LLMs that review decisions for fairness, consistency, and compliance before execution [79]. This decentralized design improves situational awareness, reduces redundant communication, and enhances system integrity through cross-agent verification.
Beyond architecture, LLM observability plays a critical role in bias mitigation by enabling inspection of intermediate reasoning steps and decision patterns in real time [35]. While observability does not eliminate bias at its source, it provides a practical mechanism for detecting and correcting biased behavior in deployed systems.
Complementing observability, Morales et al.[86] propose LangBiTe, a model-driven testing framework that enables continuous ethical assessment through domain-specific specifications and automated scenario generation. Such structured testing provides additional assurance for bias detection and accountability in LAUS.
V-C Ethical LLMs in Action: Ensuring Accountability and Transparency for UAV Operation and Communication
The deployment of LLMs in critical domains exposes a fundamental tension between capability and accountability. LLMs often function as opaque black boxes, making it difficult to interpret decisions, trace failures, or ensure consistent behavior. This opacity is exacerbated by the proprietary nature of many state-of-the-art models, limiting access to architectural and training details. Enhancing causal reasoning offers a promising pathway to improved interpretability. By modeling cause-and-effect relationships rather than relying solely on pattern recognition, LLMs can provide more transparent and auditable decision logic, enabling error diagnosis and accountability in complex, real-world UAV tasks.
This principle is exemplified by frameworks such as FlyAdapt, which integrates an LLM-assisted causal reasoning engine with predictive edge caching and hybrid verification for multi-UAV systems. By prioritizing interpretability and real-time validation, FlyAdapt achieves substantial latency reduction while intercepting the majority of diverse attacks, demonstrating how transparent, modular design can support reliable deployment in high-stakes environments. To manage residual risks, LLM observability must be adopted as a foundational practice. Observability connects input quality to output accountability through continuous monitoring, automated evaluation, and human feedback across the AI lifecycle. These mechanisms mitigate hallucination and performance drift, transforming LLMs from opaque components into governable systems.
In summary, accountable LLM deployment requires advances in intrinsic interpretability through causal reasoning, transparent system architectures embedding verification, and rigorous observability for continuous governance. Together, these approaches enable secure and trustworthy integration of LLMs into UAV systems [50].
V-D Sustainable Intelligence: Environmental Concerns in LLMs
The environmental footprint of LLMs presents a growing ethical challenge that requires a lifecycle perspective encompassing training, inference, deployment, and hardware production.
From an energy perspective, training dominates carbon emissions, with large-scale models consuming electricity comparable to hundreds of households, while inference incurs smaller per-query costs. However, at the scale of billions of daily interactions, inference energy consumption becomes a significant contributor to global emissions. Environmental impacts are further amplified by lifecycle factors such as resource-intensive hardware manufacturing, embodied carbon in AI accelerators, and regional variation in data-center energy mixes. Design trade-offs also matter: larger models may be more efficient per high-quality output, whereas smaller models can be preferable under high request volumes.
Importantly, sustainable trajectories remain achievable through mitigation strategies including model quantization, energy-efficient accelerators, deployment on low-carbon infrastructure, and standardized lifecycle environmental reporting. Aligning model design, infrastructure, and governance with transparency and renewable energy use is essential for reducing long-term ecological impact. In summary, while per-query emissions remain modest, the aggregate and systemic environmental impacts of LLMs are substantial. Achieving sustainable intelligence requires coordinated optimization across computation, energy sourcing, hardware lifecycles, and transparent environmental governance [119],[37].
V-E Key Findings and Insights
Recent studies consistently show that effective integration of LLMs into UAV systems depends not only on model capabilities but also on system-level mechanisms that ensure reliability, safety, and accountability. In practice, this calls for a layered governance framework combining LLM observability, explicit safety verification, and HITL oversight.
LLM observability forms the foundation of trustworthy deployment by linking output quality with accountability. Unlike traditional monitoring focused on latency or resource usage, observability enables continuous assessment of LLM outputs across correctness, relevance, coherence, and factual consistency, allowing early detection of model drift, hallucinations, and performance degradation in dynamic UAV environments [12]. Modern observability extends classical logs and traces with LLM-specific signals, including structured prompt logging, user feedback, request tracing across multi-stage pipelines, and token-usage monitoring, supporting systematic diagnosis and scalable optimization [89], [88].
However, observability alone is insufficient to guarantee safety. Due to their probabilistic nature, LLMs may still produce decisions that violate operational or safety constraints. A second key insight is therefore the necessity of an independent safety verification layer, where LLM-generated decisions are evaluated against predefined rules and feasibility constraints and overridden when unsafe. This separation of adaptive reasoning and deterministic enforcement enables UAV systems to combine flexibility with the reliability required for safety-critical operation. Representative observability and verification components are summarized in Table XIII.
Beyond automated safeguards, HITL oversight remains indispensable. While LLMs provide strong reasoning and decision-support capabilities, they are inherently limited in handling ethical dilemmas, rare edge cases, and ambiguous real-world scenarios. HITL frameworks integrate human judgment into mission planning and real-time operation, allowing operators to intervene when uncertainty or risk is high [108]. Effective HITL systems balance automation with oversight by reserving human intervention for complex or high-stakes decisions, supported by clear role definitions, intuitive interfaces, escalation criteria, and continuous feedback loops. Diversity among human reviewers further reduces bias and strengthens accountability [69], [26].
Taken together, these findings indicate that robust LAUS should be designed as governed socio-technical systems rather than standalone intelligent agents. LLM observability provides transparency and quality control, safety verifiers enforce hard constraints, and HITL mechanisms supply ethical judgment and resilience, enabling UAV systems to scale intelligence while preserving safety, trust, and accountability in real-world deployment.
VI ICL-Based Case Study for Multi-UAV Systems
This section transitions from theoretical frameworks to practical validation by presenting an original case study on ICL for multi-UAV coordination. We demonstrate the application of an edge-hosted LLM to jointly optimize data-collection scheduling and velocity control in a dynamic emergency-response scenario. The case study details the system architecture, simulation environment, and operational workflow of the proposed ICL-based framework. Finally, we present comparative performance results against traditional RL baselines, empirically highlighting the advantages of LLM-assisted adaptability and real-time decision-making in UAV networks.
This case study applies the ICL framework to multiple UAVs that jointly optimize data collection scheduling and velocity control. When multiple UAVs operate simultaneously, the actions of each UAV influence not only its own future states but also the behavior of other UAVs, resulting in a coupled multi-agent decision problem.
In this scenario, each UAV interacts with an unknown environment and periodically transmits collected data and state information to an LLM. The LLM uses four key components – the task description ($TD_{\text{task}}$), example set ($Ex_{t}$), current environment state ($e_{t}$), and its internal reasoning – to generate the next action ($a_{t}$). The task description encodes objectives, constraints, and expected outcomes, enabling operators to avoid explicitly designing optimization models. During operation, the LLM iteratively updates its decisions based on environment feedback, example demonstrations, and prior UAV actions, gradually converging on an effective policy.
VI-A Multi-UAV Interaction and Decision Coupling
(a) DCS-ICL vs MADQN packet loss comparison
(b) Performance comparision of different LLMs.
Figure 10: Comparative analysis of packet loss over time steps between the proposed DCS-ICL method and the baseline MADQN approach along with performance evaluation of different LLMs.
In the multi-UAV scenario, each UAV independently selects sensors and adjusts its velocity. However, these decisions are interdependent and require a multi-agent ICL formulation to jointly minimize the total system cost, accounting for factors such as packet loss, buffer overflow, and failed transmissions.
We propose the ICL-based Data Collection Scheduling and Velocity Control (DCS-ICL) framework for UASNets. DCS-ICL leverages ICL to optimize emergency data collection schedules and UAV velocities. Specifically, each UAV gathers sensory information, including queue lengths, battery levels, channel conditions, and UAV locations, from ground sensors and transmits this data to an edge-hosted LLM. The LLM is initialized with a structured system prompt that defines its operational role, constraints, and ethical safeguards, thereby ensuring compliance with safety and security requirements.
Upon receiving environmental data, the LLM generates a detailed task description in natural language that includes mission objectives, input parameters, operational rules, illustrative examples, expected outputs, and a feedback mechanism. In this context, the LLM infers an optimal data-collection schedule and velocity, which the UAVs then execute. Performance metrics and updated environmental states are continuously monitored and fed back into subsequent decision cycles, enabling dynamic adaptation and improvement of scheduling effectiveness over time.
VI-B Simulation Setup
We consider a scenario in which 10 ground sensors are randomly placed within a $100\penalty 10000\ \text{m}\times 100\penalty 10000\ \text{m}$ area. Each sensor is configured with a maximum battery capacity of $50\penalty 10000\ \text{J}$, a transmit queue limit of 40 packets, and a maximum transmit power of $100\penalty 10000\ \text{mW}$. In the learning configuration, each episode consists of 30 time steps. Simulations are performed in a customized Python 3.5 environment. UAV behavior is modeled according to the policies learned by the DCS-ICL framework. The wireless connection between UAVs and sensors is implemented using the probabilistic channel model. All experiments are conducted on a Lenovo workstation running Ubuntu 20.04 LTS, equipped with an Intel Core i5-7200U CPU @ 2.50 GHz × 4 and 16 GB of RAM. The baseline technique is Multi-agent Deep-Q-Network (MADQN): Each UAV optimizes its data collection schedule and velocity using DQN by dynamically adapting to key environmental conditions, including UAV locations, queue length, battery level, and channel conditions, to minimize packet loss.
VI-C Performance Evaluation
Fig. 10(a) illustrates the convergence of the proposed approach using GPT-4O-mini, compared to the MADQN baseline (MADQN performance is reported for the last episode). Both the proposed approach and MADQN achieve comparable performance. Fig. 10(b) illustrates the temporal evolution of packet loss for the compared models. As observed, all methods start with a relatively high packet loss at the initial time step. However, Grok and the two Mistral-based models exhibit rapid convergence, reducing packet loss to zero within the first two to three time steps and maintaining stable zero-loss performance thereafter. In contrast, LLMA shows only a partial reduction, stabilizing at a non-zero packet loss level of approximately eight packets and failing to achieve full convergence throughout the evaluation horizon. These results demonstrate the superior convergence speed and stability of the proposed approach relative to LLMA, highlighting its effectiveness in quickly mitigating packet loss and sustaining reliable network performance.
VII Most recent trends
The integration of LLMs into specialized domains, such as UAV networks, presents both unprecedented opportunities and complex challenges. Although current models demonstrate remarkable capabilities in natural language understanding, reasoning, and autonomous decision-making, limitations persist in areas such as domain-specific expertise, efficiency, interpretability, and sustainability. Addressing these gaps requires targeted research that improves model performance while ensuring practical, ethical, and environmentally conscious deployment. The following directions highlight key avenues for advancing LLMs in UAVs and related applications, including improvements in specialization, efficiency, privacy, reliability, and integration with emerging technologies such as AGI and quantum computing.
VII-A Near-Term Directions
VII-A1 Domain-Specific UAV LLMs for Autonomous Aerial Intelligence
A salient direction for future research is the development of fine-tuned, domain-specific LLMs explicitly designed for UAV applications. Drawing on industry trends in finance and healthcare, a specialized ”UAV-LLM” would be pretrained on aeronautical manuals, flight regulations, and telemetry data, then fine-tuned for mission-specific tasks such as trajectory planning, anomaly detection, and real-time decision-making. This deep specialization is expected to reduce hallucinations, improve accuracy, and provide a reliable cognitive core for UAV autonomy, enhancing operational safety, mission efficiency, and natural language interaction with ground control.
Future research should also address the environmental and economic costs associated with large-scale LLMs. Rather than indiscriminate parameter scaling, efforts should focus on domain-specific optimization, carbon accountability, and strategic investment in practical AI applications. By combining specialized model development with sustainable and cost-effective practices, UAV LLMs can deliver high performance and responsible deployment, ensuring scalable, trustworthy, and environmentally conscious aerial intelligence systems.
VII-A2 Small and Edge-Deployable LLMs for UAV Applications
The growing viability of small LLMs offers a promising research direction for UAV applications. These compact models can run efficiently on CPUs or entirely on-device, significantly reducing inference costs, enhancing privacy, and enabling deployment in offline or secure environments. Techniques such as QLoRA, model distillation, and FlashAttention have improved the performance of small models, allowing them to match or surpass larger LLMs on targeted tasks while remaining lightweight and practical for real-world workloads.
Future research should investigate the development of domain-specific, task-oriented LLMs optimized for UAV networks that achieve high accuracy, reliability, and efficient resource use. Platforms such as Google’s AI Edge Gallery demonstrate the potential for running advanced LLMs locally on devices, supporting offline execution, low latency, and privacy-preserving operations [2]. Leveraging such edge-deployable LLMs could enable real-time onboard intelligence for flight control, mission planning, and sensor data processing. Further research should also examine sustainable AI practices that balance performance with environmental impact, as well as scalable deployment strategies to maximize ROI, accessibility, and operational efficiency in UAV and other resource-constrained domains.
VII-B Medium-Term Directions
VII-B1 Self-Improving LLMs for UAV Applications
A promising direction for future research is the development of self-improving LLMs capable of generating synthetic training data to address the critical shortage of domain-specific datasets in the UAV sector. By autonomously creating and fine-tuning on curated, self-generated question-answer pairs, these models can achieve significant performance gains. In the UAV context, this approach could enable the generation of large-scale datasets covering operational scenarios, including navigation commands, anomaly detection procedures, and sensor data interpretation. Leveraging self-training capabilities enables continuous refinement of domain expertise, reduces reliance on costly manual data collection, and accelerates the development of highly specialized, efficient LLMs. Research in this area could unlock more advanced, autonomous aerial capabilities and serve as a blueprint for applying self-improving models in other data-scarce domains.
VII-B2 Reducing Hallucinations in LLMs
Addressing hallucinations in LLMs represents a critical area for future research. Hallucinations occur when models confidently generate plausible but incorrect information and can be classified as extrinsic – misrecalling facts present in training data – or intrinsic – contradicting or exceeding the input context [41]. These errors often result from ambiguous prompts, noisy or limited training data, restricted context windows, and the model’s tendency to produce answers regardless of certainty. High-stakes domains, such as telecommunications, are particularly vulnerable to the consequences of hallucinations. Future research should explore strategies to mitigate these errors, such as grounding models in verified data sources, refining prompt design, decomposing complex tasks, leveraging domain-specific models, and integrating human oversight. Advances in training methodologies, data integration, and evaluation frameworks could further enhance model reliability. Progress in these areas will be essential for developing trustworthy LLMs that consistently align with real-world knowledge and expectations.
VII-B3 Towards Reliable LLM Systems:
A clear future research direction is the development of orchestration layers and system-level frameworks for LLMs. While current LLMs are already highly capable, their limitations in hallucination, reasoning, consistency, and context handling prevent their deployment in real-world applications with full reliability. Future research should focus on designing robust systems that manage, combine, and control LLMs, rather than simply scaling model size. Key areas include: a) Agent frameworks that enable LLMs to interact with tools, APIs, and other models in structured workflows, breaking complex tasks into verifiable subtasks; b) Reliability tooling, such as automated validators, fact-checkers, and monitoring systems, to detect and correct errors before outputs reach users; c) Safety and ethical guardrails that prevent biased, harmful, or unsafe outputs through constraint mechanisms, moderation layers, and verification pipelines; d) Workflow automation that integrates HITL checkpoints and ensures repeatable, traceable, and efficient processes for high-stakes tasks.
Research in these areas will shift the focus from building ”bigger” models to creating trustworthy, production-ready LLM systems, enabling broader adoption in UAV networks while mitigating inherent limitations. The ultimate goal is to develop LLM-assisted systems that are not only fluent and capable but also reliable, safe, and self-regulating within a larger operational context.
VII-C Long-term Directions
VII-C1 AGI-Enhanced LLMs
Advancing LLMs with AGI-inspired enhancements offers a promising direction for future research. Although transformer-based LLMs have demonstrated impressive abilities in language understanding and generation, they remain limited in reasoning, adaptability, and real-world comprehension. Research aimed at integrating AGI-driven cognitive frameworks could enable LLMs to incorporate symbolic reasoning, causal inference, and dynamic knowledge updates, supporting deeper conceptual understanding beyond statistical pattern recognition. These advancements could also improve bias mitigation, ethical governance, and multimodal reasoning, ultimately allowing LLMs to generalize knowledge more effectively across the UAV domain.
VII-C2 Sustainable LLM Infrastructure
The high computational and energy demands of training and deploying LLMs remain a major constraint in the global AI landscape. For example, LLaMA-3-70B emits approximately 13 kg of $CO_{2}$ per 1 million-token batch, while GPT-3’s training consumed 1,287 MWh – enough to power 120 U.S. households for a year. Inference adds further energy load, accounting for roughly 60% of total ML energy consumption. Regional factors, such as water-intensive cooling in Arizona or fossil-fuel-heavy electricity grids in Chile, amplify environmental impacts. By 2030, AI could consume up to 10% of global electricity, exceeding the entire output of some countries, including Belgium.
VIII Conclusion
The integration of LLMs and MLLMs into UAV systems marks a significant step toward intelligent and adaptive aerial operations. This survey examined the convergence of LLM technologies with UAV communications, control, and ethical frameworks, providing a unified view of system architectures, methodologies, and real-world applications. By enabling advanced reasoning, natural language interaction, and context-aware decision-making, LLMs support applications ranging from real-time SAR to swarm coordination and mission planning. Techniques such as RAG and prompt engineering allow UAVs to operate dynamically in complex environments without extensive retraining, while MLLMs further extend these capabilities through multimodal perception and human-swarm interaction. Despite their potential, LLM-assisted UAV systems raise critical challenges related to bias, transparency, accountability, and environmental impact. Addressing these issues requires robust governance mechanisms, including HITL oversight and LLM observability, to ensure safe and trustworthy deployment. Looking forward, key research directions include domain-specific UAV LLMs, efficient edge deployment, hallucination reduction, enhanced multimodal reasoning, and sustainable AI practices. The continued synergy among LLMs, 5G/6G connectivity, edge AI, and emerging computing paradigms is expected to accelerate the evolution of intelligent UAV ecosystems, enabling more reliable and responsible autonomous aerial systems for societal applications.